### v3

# Boost Your NeRF: A Model-Agnostic Mixture of Experts Framework for High Quality and Efficient Rendering

Abstract
Since the introduction of NeRFs, considerable attention has been focused on improving their training and inference times, leading to the development of Fast-NeRFs models. Despite demonstrating impressive rendering speed and quality, the rapid convergence of such models poses challenges for further improving reconstruction quality. Common strategies to improve rendering quality involves augmenting model parameters or increasing the number of sampled points. However, these computationally intensive approaches encounter limitations in achieving significant quality enhancements. This study introduces a model-agnostic framework inspired by Sparsely-Gated Mixture of Experts to enhance rendering quality without escalating computational complexity. Our approach enables specialization in rendering different scene components by employing a mixture of experts with varying resolutions. We present a novel gate formulation designed to maximize expert capabilities and propose a resolution-based routing technique to effectively induce sparsity and decompose scenes. Our work significantly improves reconstruction quality while maintaining competitive performance. Project page: https://eidoslab.github.io/boost-your-nerf/.
1 Introduction
Neural Radiance Fields (NeRFs) [25] have recently shown impressive results in synthesizing photo-realistic 3D scenes from a set of 2D images. However, NeRFs suffer from limited scene diversity, long training time, and sensitivity to training data [22]. Since the introduction of NeRFs, significant attention has been directed toward improving their training and inference times, resulting in the development of Fast-NeRFs. By using auxiliary data structures such as voxel grids to store scene geometry and skip empty spaces, the training and inference process can be accelerated by several orders of magnitude. As a result, the neural component of these models, which is usually used to transform learned features into view-dependent color representations, becomes much smaller and can sometimes be replaced entirely by spherical harmonics [11]. Despite Fast NeRFs being an improvement from the original NeRF framework, achieving high-quality and efficient reconstructed geometry is still an open problem.
Typical “naive” approaches to enhance the reconstruction quality of these models include:
  1. 1.
    Increasing the parameters and resolution of the used data structures (e.g., voxel grid, hash grid, etc.).
  2. 2.
    Increasing the number of sampled points per ray.
  3. 3.
    Increasing the number of parameters in the neural network or the order of spherical harmonics.
However, as evident from Figure 1, the increase in reconstruction quality results in a significant increase in computational costs. In the first approach, increasing the resolution can lead to a significant improvement in the reconstruction quality, but a plateau is reached beyond which overfitting occurs, and reconstruction quality degrades. This comes at the expense of a considerable increase in spatial complexity (especially with dense voxel grids) and training times. Increasing the number of sampled points per ray can marginally improve reconstruction quality further but at a significant increase in computational complexity. In fact, the higher the number of sampled rays, the higher forward passes through a neural network are needed. Augmenting the number of parameters in the neural network is another solution. However, as the neural network grows in depth and width, it tends towards a fully implicit model, deviating from the principles of fast models (limiting the neural part as much as possible - or even removing it altogether).
Figure 1: Different strategies for improving reconstruction quality with Fast-NeRFs (DVGO [30]). Increasing the resolution of data structures like voxel grids can improve render quality, but only up to a certain point, after which quality declines (left). The MLP component’s impact on rendering quality is analyzed by varying its depth and width while keeping other variables constant (center). The effect of the number of sampled points along each ray is examined by decreasing the step size (right). In gray, our method’s performance, which significantly improves PSNR at a low computational cost.
We propose a technique capable of significantly enhancing the reconstruction quality of such models while maintaining competitive training and rendering times. Inspired by the Sparse Mixture of Experts (MoE) paradigm [29], we have developed a model-agnostic framework capable of improving the reconstruction quality of various state-of-the-art models. Intuitively, our mixture of experts consists of different-capacity (resolution) models. During training, each model specializes in rendering the most suitable parts of the scene, i.e. low-resolution models render low-frequency parts of the scene, while high-resolution models render high-frequency parts. Our formulation of the gate allows our method to be model-agnostic, enabling, on the one hand, the insertion of the gate function at the early stages of the MoE mechanism and, on the other hand, the multiplication of the gate’s output with the experts’ output as late as possible. This maximizes the capabilities of the mixture of experts and allows for working directly with the output of each expert, which is considered a black box. This is why our method is inherently model-agnostic. In summary, our contributions are as follows:
  1. 1.
    We propose the first model-agnostic framework based on Sparse MoE of models at different resolutions, which significantly improves the rendering quality of such models while maintaining competitive training and inference times (Sec. 3).
  2. 2.
    We provide a novel gate formulation inspired by Fast-NeRF models, which maximizes the capability of the mixture of experts (Sec. 3.3).
  3. 3.
    We introduce a new routing technique based on resolution, favoring the assignment of tokens to low-resolution models and discouraging the assignment of tokens to high-resolution models, inducing increasing sparsity in high-resolution models and decomposing the scene based on frequency (Sec. 3.5).
  4. 4.
    We conduct extensive experiments to test our method, including different NeRFs architectures and datasets, showing higher rendering quality and efficiency (Sec. 4).
2 Related works
2.1 Neural Radiance Field
NeRF [25] (Neural Radiance Field) has emerged as a prominent method for synthesizing novel views, showing significant progress. This approach requires a moderate number of input images along with their known camera poses. Unlike traditional methods that rely on explicit and discretized volumetric representations such as voxel grids and multiplane images, NeRF employs a coordinate-based multilayer perceptron (MLP) to create an implicit and continuous volumetric representation. NeRFs can represent a 3D scene as a MLP $F_{\theta}$, with $\theta$ being the set of its trainable parameters, such that $F_{\theta}:(\mathbf{x},\mathbf{d})\rightarrow(\mathbf{c},\sigma)$ which maps $(\mathbf{x},\mathbf{d})$, a 3D position $\mathbf{x}$ and a viewing direction $\mathbf{d}$, to a view-dependent color emission $\mathbf{c}$ and density value $\sigma$. To render the color of a pixel $\hat{C}(\mathbf{r})$, a ray $\mathbf{r}$ traverses the center of the camera through the pixel of the image plane from an origin point $\mathbf{o}$ to a ray having position $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$. Then, $N$ points are sampled on $\mathbf{r}$, and the MLP is queried for each point, obtaining a density and a color value. Finally, these results are accumulated into a single color with the volume rendering equation [23]:
$\hat{{C}}(\mathbf{r})=\sum_{i=1}^{K}T_{i}\alpha_{i}c_{i},\qquad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),\qquad\alpha_{i}=1-\exp(\sigma_{i}\delta_{i}),$ (1)
where $\alpha_{i}$ is the probability of termination at point $i$ and $T_{i}$ is the accumulated transmittance from the near plane to point $i$. NeRFs are trained by minimizing a photometric loss, which is an L2 loss between the rendered and the ground truth pixels. In more detail, given a batch $B$ of randomly sampled rays, the loss is defined as
$L_{\text{nerf}}=\frac{1}{|B|}\sum_{r\in B}\left\|\hat{C}(\mathbf{r})-C(\mathbf{r})\right\|^{2}_{2},$ (2)
where $\hat{C}(\mathbf{r})$ is the predicted color and $C(\mathbf{r})$ the ground truth color for the ray $\mathbf{r}$.
2.2 Fast Neural Radiance Fields
Since the introduction of NeRF, significant effort has been directed towards the development of faster models. Several studies focus on speeding up rendering times, for example by working on ray sampling efficiency [4, 7, 13, 38], by integrating explicit volumetric representations [12, 16, 32, 34, 33, 12, 6] or by utilizing thousands of tiny MLPs [27]. However, all these methods still require large training times. One noteworthy development is represented by the introduction of explicit volumetric representations in the training pipeline, directly optimizing such representations [11, 30, 5, 26, 10, 14, 17]. These models enable fast training and inference, with render quality only slightly inferior to full-implicit models [1, 3]. We refer to these as Fast-NeRFs. Plenoxels [11] is the first important work in this context, as it represents a scene as a sparse 3D grid where each voxel stores spherical harmonic coefficients and density. Spherical harmonics serve as an orthogonal basis for functions defined over the sphere and thus can be used for computing view-dependent color emission, without the need for a multi-layer perceptron. Additionally, Plenoxels demonstrated the advantages of linearly interpolating voxels, facilitating the learning of a continuous plenoptic function throughout the volume, akin to NeRF, albeit with discrete data. Another notable work is DVGO [30]. Each scene is there represented as two dense voxel grids (one for density and one for feature colors) alongside a MLP, for learning view-dependent color. Similar to Plenoxels, the value of each voxel is linearly interpolated with the 8 nearest voxels, but after applying the activation functions. DVGO also incorporates a preliminary coarse geometry stage to learn the scene’s general structure, facilitating adjustments to the bounding box and enabling a more intelligent ray sampling strategy. Despite utilizing lower-resolution models, DVGO achieves high reconstruction quality. The major drawback of these models is their substantial memory storage requirements. TensoRF [5] addresses this challenge by replacing the dense voxel grid with a planes and vectors decomposition, significantly reducing storage demands while maintaining comparable performance and rendering quality. Alternatively, Instant-NGP [26] proposed a multi-resolution voxel grid encoded via a hash function. They define $L$ hash grids of increased resolutions: each entry of each hash grid has $2^{T}$ parameters and $F$ features. This enables even faster training times and real-time rendering performances while maintaining a compact model; moreover, it demonstrates the efficacy of a multi-resolution approach in enhancing render quality. Inspired by this methodology, K-Planes [10] introduces a multi-resolution planar factorization of 3D space, offering reconstruction quality and model compactness similar to the previous methods, but with slightly larger training times. A similar multi-resolution planar factorization has also been proposed in Tri-MipRF [14] for mitigating the aliasing in distant or low-resolution views and blurriness in close-up shots. Given their properties, we decided to evaluate our paradigm using three different models: a 3D dense voxel grid-based model (DVGO), a decomposed grid-based model (TensoRF), and a multi-resolution hash grid-based model (Instant-NGP).
2.3 Mixture of Experts and Sparse MoE
The Mixture of Experts (MoE) paradigm [15] has gained prominence in various machine learning applications. MoE consists of multiple expert networks, each specializing in different regions of the input space, with a meta-network determining the contribution of each expert to the final prediction. Building upon the MoE framework, Sparse Mixture of Experts [29, 20, 28, 37] constitutes a scalable and efficient variant. At the core of all Sparse MoE algorithms lies an assignment problem between tokens and experts. One approach to tackle this is by approximating the solution with a gating function, which learns to assign input tokens to the most suitable experts. It typically comprises a linear layer, a softmax activation, and a Top-$K$ (where $1\leq K\leq 2$) operation, aimed at routing the input token to only a subset of the experts. To balance the assignment across all the experts, auxiliary loss functions penalizing unevenly distributed routing are often employed. This sparsity-inducing technique significantly reduces computational overhead while preserving the expressive power of the MoE architecture.
2.4 MoE and NeRF
Combining the sparse mixture of experts’ paradigms with neural radiance fields presents a non-trivial challenge. In recent research, Switch-NeRF [36] has been introduced as a novel end-to-end large-scale NeRF with learning-based scene decomposition. Inspired by Fedus et al. [9], they propose a full-implicit model with an MLP-based gate comprising 4 layers with 128 neurons and a Top-$1$ function. However, their architecture is ad-hoc, and suffers from extensive training times. This poses a challenge for us, as we aim to consider each expert as a black box, taking a point in space and a direction as input and outputting radiance and density values. Moreover, as the mixture comprises experts at different resolutions, we also aim to prioritize routing input tokens toward lower-resolution models and minimize the usage of high-resolution models.
Our work intends to overcome these limitations. We design a novel gate formulation inspired by Fast-NeRFs, that guarantees fast convergence and better performances. We position the gate at the beginning of our MoE pipeline and postpone the multiplication of the gate output with expert predictions as much as possible. This allows our framework to be entirely model-agnostic. Furthermore, our mixture of experts maintains competitive training times, enabling, in the worst case, training models on complex scenes in approximately 1 hour while achieving state-of-the-art accuracy.
3 Method
In this section, we introduce our model-agnostic ensembling approach. We present in Sec. 3.1 an overview of our method. We also describe the ray sampling and filtering mechanism in Sec. 3.2. Then, in Sec. 3.3, we discuss the effective construction of a mixture of experts with a gating mechanism in the context of NeRF. Finally, in Sec. 3.5, we present a novel formulation for decomposing the scene into resolution-based parts with an end-to-end training procedure.
3.1 Overview
Figure 2: A density field is used to compute density values for sampled points along a ray. A filtering step discards low-density points, routed through gating network $G$ for expert assignment. Top-$K$ experts compute radiance and density, aggregating and weighting these values by the corresponding gating probability to get the final values $\mathbf{c}_{f}$ and $\sigma_{f}$. Volume rendering equation yields pixel colors, and joint optimization refines our resolution-weighted auxiliary loss, allowing for high-quality and efficient rendering.
Figure 2 shows a high-level scheme of our method. After independently training a set of $M$ NeRF models $\{E_{i}\}_{i=1}^{M}$ at different resolutions, we sample $N$ points along a ray $\mathbf{r}$. For each point $\mathbf{x}_{p}$, a density value $\sigma$ is calculated, and based on this, a filtering step is performed based on the density volume $\mathbf{V}_{D}$, discarding points in regions with a density below a certain threshold $T$. Initializing the density volume $\mathbf{V}_{D}$ with the one learned from the lowest resolution model can be helpful (though it can also be learned). Subsequently, the filtered point $\mathbf{x}_{f}$ is fed to the gate $G$, so we denote with $G(\mathbf{x}_{f})$ the probabilities with which the point is assigned to the $M$ experts. Based on these values, each point is routed to the Top-$K$ experts. Each selected expert then computes the radiance $\mathbf{c}_{i,f}$ and density values $\sigma_{i,f}$ for the point, both of which are multiplied by their respective probabilities and summed together to get the final radiance $\mathbf{c}_{f}$ and density $\sigma_{f}$. Subsequently, the volume rendering Equation 1 is used to aggregate all colors and compute the pixel color for the given ray. Finally, we compute the total loss and jointly optimize the gate and the experts.
In the next sections, we will dive into the gating and sparse mixture of expert modules.
3.2 Ray Sampling and Filtering
The first phase of our method involves learning a coarse and explicit density field starting from the density volume $\mathbf{V}_{D}$, that we can leverage for skipping empty spaces. Given a ray $\mathbf{r}$ as explained in Sec. 2.1, we sample $N$ points along it:
$\mathbf{X}_{p,r}=\operatorname{sampling}(\mathbf{r})\in\mathbb{R}^{N\times 3},$ (3)
with $p\in[0,...,N-1]$. We suppress the index $r$ for abuse of notation. Next, we compute the density for each point $\mathbf{x}_{p}\in\mathbb{R}^{3}$ and discard those with negligible density. The density value $\mathbf{\sigma}_{D,p}$ for each point is computed by linearly interpolating $v$ neighboring voxels:
$\sigma_{D,p}=\operatorname{act}(\operatorname{interpolate}(\mathbf{x}_{p},V_{D}))\in\mathbb{R}$ (4)
where $\operatorname{act}$ represents a density activation function (such as softplus). We denote with $\mathbf{x_{f}\in\mathbb{missing}}{R}^{3}$ a remaining point after the filtering operation.
3.3 Trainable Gating Model
Our gating mechanism incorporates a hybrid architecture: an explicit feature grid $\mathbf{V}_{G}$ and a shallow MLP. The gating mechanism can also be seen as a probability field. First, we compute per-point features:
$\mathbf{feat}=\operatorname{interpolate}(\mathbf{x}_{f},\mathbf{V}_{G})\in\mathbb{R}^{C}$ (5)
where $C$ represents the number of channels of $\mathbf{V}_{G}$. Subsequently, we transform each feature into a probability. We compute logits as
$\mathbf{logits}=\operatorname{MLP}(\mathbf{feat})\in\mathbb{R}^{M}$ (6)
where $M$ is the number of experts in the mixture of experts. Finally, we apply a per-row softmax. Our gating function can be summarized as:
$G(\mathbf{\mathbf{x}_{f}})=\operatorname{softmax}(\operatorname{MLP}(\operatorname{interpolate}(\mathbf{x}_{f},\mathbf{V}_{G})))\in\mathbb{R}^{M}$ (7)
3.4 NeRFs Experts
As mentioned in Sec. 3.1, we employ a set of $M$ pretrained NeRFs models as experts. After feeding as input each point $\mathbf{x}_{f}$ to the gate, we route them to the $k$ experts with the $k$ highest probabilities $G(\mathbf{x}_{f})$. We define this set of indexes of the selected experts as
$\mathcal{K}=\operatorname*{argmax}_{Top-k}(G(\mathbf{x}_{f})).$ (8)
Each expert is a Fast-NeRF model that receives the dispatched points and the direction of the ray they lie on as input and outputs a density and a radiance value. We treat each expert as a black box, ensuring our method is inherently model-agnostic. The radiance $\mathbf{c}_{f}$ and the density $\sigma_{f}$ for the point $\mathbf{x}_{f}$ laying on a ray with direction $\mathbf{d}$ can be expressed as the sum of each expert predictions weighted by their probability:
$\mathbf{c}_{f},\sigma_{f}=\sum_{i\in\mathcal{K}}E_{i}(\mathbf{x_{f}},\mathbf{d})\cdot G(\mathbf{x_{f}})_{i}$ (9)
where $E_{i}$ denotes the $i$-th expert and $G(\mathbf{x_{f}})_{i}$ the probability for the point to be dispatched to that expert $i$.
Once we obtain radiance and density values for all the points on the ray, we can compute the pixel color with Equation 1.
3.5 Resolution-based Routing
To balance the load and prevent the gate from focusing on assigning points to a single expert (typically the one with the highest resolution), we employ a resolution-weighted auxiliary loss. Given $M$ experts, and a batch of points $B$, the auxiliary loss is defined as:
$L_{\text{aux}}=\frac{M}{|B|^{2}}\sum_{i=1}^{M}c_{i}m_{i},\qquad m_{i}=\sum_{\mathbf{x}_{f}\in B}G(\mathbf{x}_{f})_{i},$ (10)
where $c_{i}$ represents the number of inputs dispatched to the expert $i$, and $m_{i}$ is the sum of all the probabilities for each point in the batch for the expert $i$. This loss helps balance the workload, ensuring that each expert processes a similar number of points. Ideally, in a perfectly balanced scenario, $L_{aux}$ is expected to be 1, as both $c_{i}$ and $m_{i}$ would be $\frac{M}{N}$, resulting in $\sum_{i=1}^{M}c_{i}m_{i}$ being $\frac{M^{2}}{N}$. However, in the case of total imbalance, it tends to the number $M$. We aim to take a step further by assigning as many points as possible to low-resolution experts while discouraging point assignment to high-resolution models. Hence, we introduce some penalty terms $w_{i}$ associated with each expert. The higher the resolution of the model, the higher the penalty. We define a novel resolution-based auxiliary loss:
$L_{\text{rw-aux}}=\frac{M}{|B|^{2}}\sum_{i=1}^{M}c_{i}m_{i}w_{i}.\qquad$ (11)
As for the weighting strategy, we opted for the following geometric progression. The weight $w_{i}$ for the $i$ expert is computed as:
$w_{i}=\exp\left(\frac{\ln M}{M-1}\right)^{i},\quad i\in[0,...,M-1].$ (12)
The total loss is then defined as:
$L_{\text{tot}}=L_{\text{nerf}}+\lambda L_{\text{rw-aux}}.$ (13)
As we will see in Sec. 4, our proposed loss not only further improves the reconstruction quality but also encourages sparsity in higher-resolution models. A comparison with other weighting strategies is proposed in the Supplementary 9.
4 Experiments
4.1 Setting
The code was written in Python 3.8, using PyTorch 1.12, and executed on a single NVIDIA A40 GPU. We tested our technique on the DVGO, TensoRF, and Instant-NGP architecture. The code from official repositories was used as the starting point. For Instant-NGP, the native implementation in PyTorch ngp_pl was employed, which exhibits comparable performance and reconstruction quality to the official NVIDIA implementation. We experimented with various configurations, ranging from a minimum of 3 models to 5 models of different resolutions. The experts are ordered by resolution, such that the $i+1$-th expert has approximately double the parameters of the $i$-th expert. For all configurations and models, $\lambda$ was set to $10^{-3}$. The Gate consists of a grid (dense voxel grid for DVGO, factorized grid for TensoRF, and hash grid for Instant-NGP) and a shallow MLP (2 layers with 64 neurons each) with ReLU activation function. The resolution of the gate is low: $128^{3}$ for both DVGO and TensoRF, while for Instant-NGP we used $L=6$. The number of iterations is set to $20k$ for all the architectures. All other hyperparameters are left as the original implementations. We draw experiments with our method using Top-$k$ experts, with $k=1$ and $k=2$. We compare our results versus baselines with comparable resolutions, as well as a Fast-NeRF ensemble (Ens) obtained by jointly fine-tuning all models and averaging their predictions. It can be noted that this ensemble can be interpreted as a limit case for our method with $k=M$ and $G(\mathbf{x}_{f})_{i}=1/M,\forall i$ similar to the method proposed by [8].
Table 1: Results on DVGO, TensoRF and Instant-NGP with $M=5$. $\|w\|_{0}$ is expressed as multiple of $10^{6}$, while for Instant-NGP as a multiple of $10^{5}$.
Dataset Metrics DVGO TensoRF Instant-NGP
baseline Top-1 Top-2 Ens baseline Top-1 Top-2 Ens baseline Top-1 Top-2 Ens
Blender $\mathbf{PSNR}\uparrow$ 33.04 33.43 33.74 33.79 32.98 33.68 34.09 34.00 33.35 33.56 33.83 34.01
$\mathbf{SSIM}\uparrow$ 0.961 0.964 0.965 0.966 0.958 0.965 0.968 0.968 0.963 0.963 0.965 0.966
$\mathbf{LPIPS}\downarrow$ 0.026 0.024 0.022 0.022 0.029 0.023 0.021 0.022 0.025 0.045 0.043 0.042
$\|w\|_{0}\downarrow$ 99 26 39 97 40 24 33 49 31 17 21 26
$\mathbf{GFLOPs}\downarrow$ 635 499 940 2206 886 732 1344 2214 46 71 123 208
$\mathbf{Time}\downarrow$ 26’ 21’ 25’ 32’ 44’ 69’ 76’ 70’ 11’ 32’ 34’ 37’
NSVF $\mathbf{PSNR}\uparrow$ 35.21 37.12 37.59 37.68 36.70 37.40 37.98 38.08 36.44 36.59 37.04 37.31
$\mathbf{SSIM}\uparrow$ 0.977 0.984 0.986 0.986 0.981 0.984 0.986 0.987 0.983 0.983 0.984 0.985
$\mathbf{LPIPS}\downarrow$ 0.015 0.009 0.008 0.007 0.013 0.009 0.008 0.008 0.010 0.023 0.022 0.020
$\|w\|_{0}\downarrow$ 95 27 43 100 42 28 38 54 30 17 20 27
$\mathbf{GFLOPs}\downarrow$ 564 430 811 1903 706 575 1053 1763 28 52 72 123
$\mathbf{Time}\downarrow$ 12’ 22’ 25’ 30’ 46’ 75’ 75’ 75’ 10’ 31’ 33’ 34’
TaT $\mathbf{PSNR}\uparrow$ 28.93 29.14 29.27 29.37 28.44 28.78 29.14 29.11 29.07 29.16 29.32 29.38
$\mathbf{SSIM}\uparrow$ 0.927 0.925 0.929 0.932 0.905 0.924 0.929 0.928 0.924 0.927 0.929 0.930
$\mathbf{LPIPS}\downarrow$ 0.107 0.108 0.105 0.103 0 0.106 0.099 0.109 0.101 0.125 0.124 0.121
$\|w\|_{0}\downarrow$ 74 16 26 65 7 9 15 80 88 37 47 70
$\mathbf{GFLOPs}\downarrow$ 2666 1626 3066 7198 3567 2791 5126 9146 211 229 531 1003
$\mathbf{Time}\downarrow$ 22’ 25’ 28’ 38’ 72’ 70’ 78’ 101’ 14’ 37’ 39’ 45’
LLFF $\mathbf{PSNR}\uparrow$ 26.24 26.43 26.62 26.65 26.71 26.73 27.09 27.10 24.97 24.90 25.17 25.19
$\mathbf{SSIM}\uparrow$ 0.831 0.832 0.839 0.843 0.835 0.836 0.862 0.864 0.764 0.763 0.777 0.778
$\mathbf{LPIPS}\downarrow$ 0.136 0.115 0.111 0.107 0.114 0.111 0.101 0.101 0.128 0.239 0.237 0.234
$\|w\|_{0}\downarrow$ 62 26 40 113 19 11 16 23 152 68 75 148
$\mathbf{GFLOPs}\downarrow$ 1678 1514 2508 4972 4542 3226 5921 13522 573 887 1391 2712
$\mathbf{Time}\downarrow$ 24’ 28’ 32’ 36’ 58’ 49’ 57’ 68’ 24’ 25’ 46’ 42’
4.2 Metrics and Datasets
For each test, image-quality metrics such as PSNR, SSIM [31], and LPIPS [35] (computed on AlexNet [19]) are presented. Additionally, we report the number of non-zero parameters as $\|w_{0}\|$, the average GFLOPs required by each model to render the images of the test set, and total training times. We present results obtained across four major datasets, namely: Synthetic-NeRF [25], Neural Sparse Voxel Field Dataset (NSVF) [21], TanksAndTemple [18], and Local Light Field Fusion Dataset (LLFF) [24].
4.3 Quantitative Results
The main experimental results in terms of the metrics defined above are shown in Table 1. Here, we present the results obtained using a mixture of $M=5$ experts; further configurations can be found in the supplementary Sec. 12. Our experiments reveal that our MoE provides a significant rendering quality improvement with respect to the baseline with no or limited impact in terms of computational cost. The increase in rendering quality is notable in Synthetic NeRF (up to 1 dB) and NSVF (up to 1.3 dB). Although more moderate, improvements are still evident, even in more challenging datasets such as TanksAndTemple and LLFF, with about 0.5 dB gain. While Top-$1$ can already achieve state-of-the-art render quality, the Top-$2$ strategy further enhances image quality, reaching levels comparable to the ensemble but with much greater efficiency (about half the average FLOPs per rendering) and using significantly fewer parameters. This observation is also illustrated in Figure 3 , where FLOPS/PSNR plots are shown (each marker in every curve refers to the cases with $M=3,4,5$ experts respectively). Training times are on average are longer but still acceptable (in the worst-case scenario, around 1h is required to train our MoE). However, in the case of Top-1, they can be faster than baselines. This is because of our resolution methods, which tend to favor low-resolution models. Based on these analyses, the Top-$2$ strategy strikes an excellent quality/cost trade-off. The ensembling configuration, using all experts, represents an upper bound in terms of image quality while being the worst case in computational terms.
(a) DVGO
(b) TensoRF
(c) Instant-NGP
Figure 3: PSNR/GFLOPs plots for Synthetic-NeRF and TanksAndTemple. The remaining datasets show similar results.
4.4 Qualitative Results
Figure 4 shows visual comparisons among the baseline, ensemble, Top-$1$, and Top-$2$.
Figure 4: Qualitative results on some of the scenes of each dataset. From left to right: ground truths, baselines, Top-1, Top-2 and Ensemble. Each model has the same parameters and has been trained for the same number of iterations (DVGO).
From the selected image crops, one can appreciate that our method ensures superior reconstruction quality compared to the baselines, effectively reproducing sharper details while reducing noise on texture-less spots. This is particularly notable when examining elements such as the window decorations in the Palace scene or the text on the box above the desk in the Room scene. Additionally, surfaces such as the semi-transparent glass in the Wineholder or the Caterpillar scene appear less noisy and more faithful to the original. This improvement can be attributed to the synergy among different resolution models: low-resolution models excel at representing lower frequencies, thereby introducing less noise, while high-resolution models can focus solely on high-frequency components. Additionally, it is important to notice that there is no significant difference between the ensemble and Top-$2$.
4.5 Comparison with Naive Methods
As shown in Figure 1, our method leads to significantly higher quality reconstructions at a greatly reduced computational cost. Decreasing the step size yields the least noticeable improvement while incurring a substantial computational expense, as each sampled point requires an evaluation by the color decoder. Our method achieves high-quality reconstructions with the same step size as the baseline models. Increasing the MLP parameters can boost accuracy, but again, the computational costs become substantial compared to our method. Similarly, increasing resolution results in way inferior PSNR with respect to our method at a comparable computational cost. Our method also allows scaling to higher resolutions, while baseline models tend to introduce noise and artifacts, leading to a decrease in reconstruction quality as the resolution increases.
4.6 Gate Visualization and Scene Decomposition
In Figure 5, we visualize the gate (probability field in grayscale) and each expert output in the case $M=3$. On the right side of the figure, one can appreciate per-model renders and experts’ specialization. It is worth noting how higher-resolution experts render high-frequency details. This is particularly evident in the Mic scene.
Figure 5: Gate (gray-scale images) and per-model output visualizations with 3 experts and resolution-based routing (DVGO). Images are ordered by increasing resolutions.
4.7 Ablation
Res Gate Type $\mathbf{PSNR}$ $\|w\|_{0}$ $\mathbf{GFLOPs}$
$128^{3}$ Ours 36.80 26 476
$256^{3}$ Ours 36.77 33 601
$300^{3}$ Ours 36.79 42 677
- Linear 36.25 24 444
- Switch-NeRF 36.44 25 1095
Table 2: Comparison of different gate resolutions and formulations on Lego scene and DVGO.
Figure 6: Comparison of Gating Functions (Top-1 to Top-5) on Synthetic-NeRF with DVGO.
Here we investigate how different design choices of our model affect the performance and rendering quality. These choices include the gate resolution, different gate formulations, and the number of experts in the Top-$k$.
Here we investigate how different design choices of our model affect the performance and rendering quality. These choices include the gate resolution, different gate formulations, and the number of experts in the Top-$k$.
4.7.1 Gate Resolution and Gate Formulation.
We show in Table 2 that a low-resolution gate is sufficient to achieve high-quality rendering. Interestingly, as the gate resolution increases, there is a slight decrease in rendering quality during testing, coupled with an increase in required FLOPs per image rendering and number of parameters. We also compare various gating strategies, including linear gating, the configuration proposed by Switch-NeRF, and our gate formulation. Through experimentation, it becomes evident that our approach outperforms others in terms of both render quality and performance, achieving performances comparable to a linear gating function.
4.7.2 Why Top-2?
Here we investigate the performance and rendering quality trends with varying values of $k$. In Figure 6, we can see that while $k=1$ can lead to a significant increase in quality with comparable performance to baseline models, the Top-2 further enhances quality at the expense of increased (but still acceptable) computation. The Top-3, Top-4, and Top-5 provide marginal quality improvements at a significantly higher computational cost. Hence, we consider the Top-2 the optimal balance between performance and quality.
5 Limitations
Our study present some limitations. First of all, a pre-training phase, on which each model is trained independently is required for achieving good results. Training end-to-end, without a pre-training phase, can lead to reconstruction quality that is noticeably inferior to baselines. Pre-training models at different resolutions allows for diversified architectures, making it easier for the gate to learn the decomposition. The training is also strongly dependent on the auxiliary loss. Different values of $\lambda$ can significantly influence performances and load balancing. We conducted a sweep to identify a value that works well across many scenarios, but it may not be suitable for different datasets. Another limitation is the overhead introduced by the MoE. Each input token is first interpolated with the gate’s grid, decoded into probabilities and routed to the chosen expert. These operations can significantly slow-down the rendering process, leading to higher training and inference times. Although considerable effort was devoted to developing the most efficient gate possible, our MoE is still slower than the respective baselines.
6 Conclusions
In this paper, we introduced a model-agnostic framework for enhancing the rendering of Fast-NeRF models. Our formulation of the Gate reduces computational costs in both training and inference phases while ensuring better quality compared to a traditional gate. Additionally, the introduction of an auxiliary loss with res penalty allows for increased utilization of low-resolution models, reducing the number of active parameters and promoting sparsity in higher-resolution models. Our results demonstrate how this approach can significantly improve reconstruction quality while considering performance metrics. Specifically, we show that a Top-2 strategy strikes a good balance between performance and quality.
Acknowledgements
This work was partially funded by Hi!PARIS Center on Data Analytics and Artificial Intelligence.
7 Additional Details
Our MoE is reducible to a Sparsely-Gated Mixture of Experts with full capacity. Since the number of points sampled is reasonable, we did not find necessary to impose a limit to the capacity factor of each model. We also introduce different-resolution experts in the MoE, promoting the routing of tokens towards lower-resolution experts. Regarding DVGO and TensoRF, we utilized models at various grid resolutions, while keeping all other hyperparameters consistent with the original implementations. The selected resolutions are $160^{3},200^{3},256^{3},320^{3},$ and $384^{3}$. For Instant-NGP, we trained models while varying the $L$ parameter across the following values: $6,8,10,16,$ and $24$. Experts within the MoE are arranged in increasing resolution order.
7.1 $\|w\|_{0}$ Computation
We define $\|w\|_{0}$ as the count of non-null parameters in the Mixture of Experts. As our MoE operates as a Sparsely-Gated MoE, only a subset of the Neural Radiance Fields models are effectively employed during rendering. Moreover, our resolution-based routing mechanism has the potential to significantly reduce this parameter.
To compute the parameters, we freeze the MoE and render all the images in the test set. Subsequently, we calculate the loss in the usual manner and compute gradients with respect to the MoE’s parameters. $||w||_{0}$ then represents the count of parameters with non-null gradients:
$\|w\|_{0}:=\|\frac{\partial L_{\text{nerf}}}{\partial W_{\text{MoE}}\neq 0}\|_{0}$ (14)
This can be associated with gradient-magnitude pruning. We leave this idea as a future project.
7.2 Instant-NGP Implementation
The implementation for Instant-NGP closely resembles that proposed in Figure 2, albeit with a slight modification: the gate comprises a tiny Instant-NGP model. This gate model consists of a multi-resolution hash grid with $L=6$, which computes per-point features. These features are then linearly interpolated with the 8 nearest voxels, similar to DVGO and TensoRF, before being concatenated together. Following this concatenation, they undergo transformation into probabilities using the same methodology as in the main method. Additionally, the MLP remains shallow, comprising 2 layers with 64 neurons each, activated by ReLU functions.
7.3 Ensembling NeRFs’ Outputs
Figure 7: Upper bound for $\Delta E_{12}$ such that (20) is satisfied.
Figure 8: Estimated performance gain as a function of $\alpha_{1}$ and of the difference in output between the two NeRFs.
To validate our Mixture of Experts, we conducted a comparison with an ensemble of Neural Radiance Fields, where the output is determined by averaging the predictions of each NeRF model. Starting from a set of $M$ pre-trained models, an ensemble is built and is then fine-tuned. Averaging the outputs of NeRFs can notably enhance the quality of reconstruction. Let us assume we have two NeRFs, having as output respectively $\boldsymbol{y}_{1}$ and $\boldsymbol{y}_{2}$, while the ground truth is $\hat{\boldsymbol{y}}$. We can write an equality of their performance, in terms of mean-squared error:
$E_{1}$ $=\Delta E_{12}+E_{2}$
$\frac{1}{M}\sum_{j}(y_{1j}-\hat{y})^{2}$ $=\Delta E_{12}+\frac{1}{M}\sum_{j}(y_{2j}-\hat{y})^{2}$ (15)
where $\Delta E_{12}$ is the error gap between the two models and $M$ is the number of outputs on which the error is averaged. From (15) we can easily write
$\frac{1}{M}\sum_{j}(y_{2j}^{2}-y_{1j}^{2})=\Delta E_{12}+\frac{2}{M}\sum_{j}\hat{y}(y_{2j}-y_{1j}).$ (16)
Through ensembling, we desire the error $E_{\text{ens}}$ to be lower than either of the two other models. Without loss of generality, let us consider the case such that it is lower than the error of the first model:
$E^{\text{ens}}$ $<E_{1}$
$\frac{1}{M}\sum_{j}(\alpha_{1}y_{1j}+\alpha_{2}y_{2j}-\hat{y})^{2}$ $<\frac{1}{M}\sum_{j}(y_{1j}-\hat{y})^{2}$ (17)
where $\alpha_{1},\!\alpha_{2}$ are two weighting factors for the two NeRF’s outputs. If we expand (17) we obtain
$\frac{1}{M}\sum_{j}(\alpha_{1}y_{1j}+$ $\alpha_{2}y_{2j})^{2}-y_{1j}^{2}<$
$\frac{2}{M}\sum_{j}\hat{y}\left[\alpha_{2}y_{2j}-(1-\alpha_{1})y_{1j}\right].$ (18)
In order to plug (15) in (18), we need to impose $\alpha_{2}=1-\alpha_{1}$. Hence, we have
$\frac{1}{M}\sum_{j}$ $\left[\alpha_{1}y_{1j}+(1-\alpha_{1})y_{2j}\right]^{2}-y_{1j}^{2}<$
$(1-\alpha_{1})\left[\Delta E_{12}+\frac{1}{M}\sum_{j}(y_{2j}-\hat{y})^{2}\right].$ (19)
By expanding and simplifying (19), we obtain
$\frac{1}{M}\sum_{j}(\alpha_{1}^{2}-\alpha_{1})(y_{1j}-y_{2j})^{2}<(1-\alpha_{1})\Delta E_{12}.$ (20)
Fig. 7 pictures the upper bound dictated by $\Delta E_{12}$ such that (20) is verified. Given that for (15) we never stated which of the two models has the lowest error, evidently if $\Delta E_{12}>0$, for $\alpha_{1}<1$ (20) is always verified. However, this is not anymore the case when $\Delta E_{12}<0$. In order to maximize the performance of the ensemble, we need to settle to the critical point of (20) (global maximum) with respect to $\alpha_{1}$, which is
$\alpha_{1}=\frac{1-\Delta E_{12}}{2}.$ (21)
If we assume $\Delta E_{12}$ being distributed as a random variable with zero means (this is a realistic assumption as we never stated which of the two models performs the best), evidently the best solution is for $\alpha_{1}=\frac{1}{2}$: Fig. 8 graphically evidences this.
8 Algorithm
We present an intuitive pseudo-algorithm outlining our Mixture of Experts framework. Our native implementation in PyTorch closely adheres to this algorithm.
Algorithm 1 Sparse MoE Trainining
1:$D$: dataset, $res$: resolution, $l$: number of pre-training iterations, $m$: number of training iterations, $M$: number of experts in the mixture, $k$: number of top points to select, $\lambda$: resolution-based aux-loss penalties, $W$: aux-loss weights
2:$experts\leftarrow$ create_experts($res$)
3:$experts\leftarrow$ pre_train($experts$, $l$)
4:$gate\leftarrow$ create_gate()
5:$moe\leftarrow$ build_moe($experts$, $gate$)
6:for $i\leftarrow 1$ to $m$ do
7:     $\mathbf{o}$, $\mathbf{d}$, $\mathbf{rgb}\leftarrow$ batch($D$)
8:     $\mathbf{x}\leftarrow$ sample_and_filter_points($\mathbf{o}$, $\mathbf{d}$)
9:     $G(\mathbf{x})\leftarrow$ $gate(\mathbf{x})$
10:     $topk\_idx$, $topk\_vals\leftarrow$ top-k($G(\mathbf{x}),k$)
11:     $\mathbf{\sigma_{s}}$, $\mathbf{c_{s}}$leftarrow array()
12:     for $j\leftarrow 1$ to $M$ do
13:         $mask\leftarrow topk\_idx==j$
14:         $\mathbf{\sigma}$, $\mathbf{c}\leftarrow$ $experts[j](\mathbf{x}[mask],\mathbf{d})$
15:         $\mathbf{\sigma_{s}}[j]\leftarrow\sigma\cdot topk\_vals[mask]$
16:         ${\mathbf{c_{s}}[j]}\leftarrow\mathbf{c}\cdot topk\_vals[mask]$
17:     end for
18:     $\mathbf{\sigma_{f}}\leftarrow$ sum($\mathbf{\sigma}_{s}$)
19:     $\mathbf{c_{f}}\leftarrow$ sum($\mathbf{c_{s}}$)
20:     $\mathbf{colors}\leftarrow$ volume_rendering($\mathbf{\sigma_{f}}$, $\mathbf{c_{f}}$)
21:     $L_{\text{tot}}\leftarrow L_{\text{nerf}}(rgb,colors)+\lambda L_{\text{rw-aux}}(P(\mathbf{x}),W)$
22:     optimize($moe$)
23:end for
The initial phase involves independently training $M$ models at different resolutions for $m$ iterations. Given that we utilize Fast NeRFs, this phase tends to be rapid. Following this, we construct the Sparsely-Gated Mixture of Experts as outlined in lines 1-4 of the algorithm. The training of the MoE commences at line 4: Given a batch of rays, typically represented by a triplet (origin, direction, and ground truth color), points along the ray are sampled and appropriately filtered. Subsequently, for each point, a probability distribution is computed, indicating the confidence of the gate in assigning each point to each expert (line 8). Based on this probability, the points are delivered to the top-k experts, who calculate, for each point, the density and radiance, which are then multiplied by their corresponding probability value (lines 9-16). The density and radiance values are aggregated, yielding a single array of density and radiance (line 20). Finally, we proceed as usual by computing the pixel color using the volume rendering equation, calculating the loss, and optimizing both the gate and the experts. Our algorithm is tailored for a single GPU, but it can be readily extended to a multi-GPU environment, as the computation of the experts’ output can be parallelized.
9 Resolution Penalties
We experimented with different resolution-based penalties, namely: linear, geometric progression, and quadratic, defined as follows:
$\text{linear}:w_{i+1}=w_{i}+k$ (22)
[…]
$\text{quadratic}:w_{i+1}=2\cdot w_{i}$ (24)
where $i\in[0,...,M-1]$ and $k=1$. Considering the results (Tab. 3), we opted for the geometric progression as it provides slightly better results. It’s worth noting how introducing such penalties further improves render quality compared to a standard auxiliary loss without resolution penalty (indicated as "none").
In Fig. 9, we provide a qualitative evaluation of experts’ specialization. With our strategy, we aim to utilize high-resolution models as sparingly as possible.
Table 3: Comparison among different penalty strategies
Strategy Top-1 Top-2
PSNR $\|w\|_{0}$ PSNR $\|w\|_{0}$
none 33.41 32 33.61 56
geom. progr. 33.43 26 33.74 39
quadratic 33.31 20 33.64 37
linear 33.41 32 33.63 36
Figure 9: Per-experts outputs with a MoE of 5 NeRFs and top-2 with different resolution penalties. Experts’ output are ordered by increasing resolution.
10 Visualizing Experts Specialization
In Fig. 10 and 11 we show the output for each expert in the MoE with $M=5$, with respectively Top-$1$ and Top-$2$ function and resolution penalty.
Figure 10: Per-expert output with Top-1 and 5 experts. Outputs are ordered from left to right by increasing experts’ resolution.
Figure 11: Per-expert output with Top-2 and 5 experts. Outputs are ordered from left to right by increasing experts’ resolution.
11 Comparison with SoTA Methods
Although our goal was to focus on fast models and improve reconstruction quality at a lower computational cost compared to their baseline counterparts, our method achieves comparable (and often superior) accuracy to explicit models. Table 4 presents a comparison with three state-of-the-art implicit models: Mip-NeRF [1], Mip-NeRF 360 [2] and Zip-NeRF [3]. Training times refers to a single GPU configuration (NVIDIA A40). Mip-NeRF was trained for $1000000$ iterations; Mip-NeRF 360 and Zip-NeRF for $200000$ iterations.
Table 4: Comparison among boosted Fast-NeRF and implicit SoTA methods on Synthetic-NeRF dataset.
[…]
DVGO Top-1 33.43 0.964 21’
TensoRF Top-1 33.68 0.965 69’
iNGP Top-1 33.56 0.963 32’
DVGO Top-2 33.74 0.965 25’
TensoRF Top-2 34.09 0.968 76’
iNGP Top-2 33.83 0.965 34’
Mip-NeRF 33.09 0.961 16h
Mip-NeRF360 32.96 0.960 5h
Zip-NeRF 33.69 0.973 7h
12 Additional Results
In this section, we present additional results for DVGO, TensoRF, and Instant-NGP with varying values of $M={3,4,5}$. In Table 5 and Figure 12, we provide aggregated results for all the tested datasets on DVGO, including plots of PSNR/GLOPS and PSNR/$\|w\|_{0}$. Similarly, in Table 6 and Figure 13, we present the results for TensoRF. The results for Instant-NGP are shown in Table 7 and Figure 14.
Table 5: Aggregated results on DVGO
Dataset Metrics M=3 M=4 M=5
baseline Top-1 Top-2 Ens baseline Top-1 Top-2 Ens baseline Top-1 Top-2 Ens
[…]
SSIM $\uparrow$ 0.962 0.962 0.964 0.964 0.963 0.963 0.965 0.965 0.963 0.964 0.965 0.966
LPIPS $\downarrow$ 0.026 0.026 0.025 0.025 0.025 0.025 0.024 0.023 0.026 0.024 0.022 0.022
$\|w\|_{0}\downarrow$ 29 17 25 32 57 20 33 59 99 26 39 97
GFLOPs $\downarrow$ 382 342 647 914 508 431 813 1529 635 499 940 2206
time $\downarrow$ 10’ 12’ 15’ 15’ 13’ 16’ 19’ 23’ 26’ 20’ 24’ 32’
[…]
SSIM $\uparrow$ 0.979 0.982 0.984 0.984 0.979 0.983 0.985 0.985 0.977 0.984 0.986 0.986
LPIPS $\downarrow$ 0.014 0.010 0.009 0.009 0.014 0.010 0.008 0.008 0.015 0.009 0.008 0.007
$\|w\|_{0}\downarrow$ 29 18 23 32 60 23 32 60 95 27 43 100
GFLOPs $\downarrow$ 318 266 502 709 452 358 676 1271 564 430 811 1903
time $\downarrow$ 7’ 13’ 16’ 16’ 22’ 17’ 21’ 25’ 12’ 21’ 25’ 35’
[…]
SSIM $\uparrow$ 0.919 0.921 0.925 0.926 0.921 0.923 0.927 0.930 0.927 0.925 0.929 0.932
LPIPS $\downarrow$ 0.131 0.115 0.113 0.113 0.121 0.110 0.108 0.107 0.107 0.108 0.105 0.103
$\|w\|_{0}\downarrow$ 17 11 18 25 30 15 22 43 74 16 26 65
GFLOPs $\downarrow$ 1299 1223 2310 3263 1547 1463 2761 5193 2666 1626 3066 7198
time $\downarrow$ 10’ 15’ 18’ 18’ 18’ 19’ 23’ 27’ 22’ 25’ 28’ 38’
[…]
SSIM $\uparrow$ 0.832 0.823 0.831 0.834 0.832 0.830 0.837 0.841 0.831 0.832 0.839 0.843
LPIPS $\downarrow$ 0.141 0.130 0.120 0.123 0.137 0.117 0.116 0.112 0.136 0.115 0.111 0.107
$\|w\|_{0}\downarrow$ 19 12 23 28 33 19 29 59 62 26 40 113
GFLOPs $\downarrow$ 976 844 1408 1693 1250 1162 1931 3079 1678 1514 2508 4972
time $\downarrow$ 12’ 14’ 16’ 13’ 16’ 20’ 23’ 23’ 24’ 28’ 32’ 36’
Figure 12: PSNR/GFLOPs and PSNR/$||w||_{0}$ plots for DVGO aggregated results.
Table 6: Aggregated results on TensoRF
Dataset Metrics M=3 M=4 M=5
baseline Top-1 Top-2 Ens baseline Top-1 Top-2 Ens baseline Top-1 Top-2 Ens
[…]
SSIM $\uparrow$ 0.962 0.965 0.967 0.966 0.962 0.966 0.968 0.967 0.958 0.965 0.968 0.968
LPIPS $\downarrow$ 0.027 0.025 0.024 0.024 0.027 0.024 0.022 0.023 0.029 0.024 0.021 0.022
$\|w\|_{0}\downarrow$ 15 11 15 18 26 16 23 31 40 24 33 49
GFLOPs $\downarrow$ 600 520 958 1327 753 641 1179 1770 886 732 1344 2214
time $\downarrow$ 29’ 40’ 44’ 40’ 36’ 55’ 61’ 53’ 44’ 69’ 76’ 70’
[…]
SSIM $\uparrow$ 0.982 0.983 0.985 0.985 0.981 0.983 0.986 0.986 0.981 0.984 0.986 0.987
LPIPS $\downarrow$ 0.011 0.010 0.009 0.009 0.012 0.010 0.008 0.008 0.013 0.009 0.008 0.008
$\|w\|_{0}\downarrow$ 17 10 15 20 28 19 26 35 42 28 38 54
GFLOPs $\downarrow$ 477 414 761 1055 591 504 924 1408 706 575 1053 1763
time $\downarrow$ 29’ 45’ 44’ 40’ 36’ 58’ 60’ 56’ 46’ 75’ 75’ 75’
[…]
SSIM $\uparrow$ 0.906 0.922 0.924 0.925 0.906 0.924 0.927 0.927 0.905 0.924 0.929 0.928
LPIPS $\downarrow$ 0.136 0.112 0.111 0.113 0.130 0.106 0.104 0.113 0.123 0.106 0.099 0.109
$\|w\|_{0}\downarrow$ 3 5 8 11 5 7 11 49 7 9 15 80
GFLOPs $\downarrow$ 2379 2151 3959 5481 2915 2475 4549 7311 3567 2791 5126 9146
time $\downarrow$ 40’ 41’ 43’ 44’ 54’ 53’ 59’ 72’ 72’ 70’ 78’ 101’
[…]
SSIM $\uparrow$ 0.834 0.835 0.841 0.840 0.836 0.839 0.843 0.843 0.835 0.836 0.862 0.864
LPIPS $\downarrow$ 0.127 0.117 0.110 0.116 0.118 0.112 0.105 0.107 0.114 0.111 0.101 0.101
$\|w\|_{0}\downarrow$ 9 5 6 8 13 6 10 14 19 11 16 23
GFLOPs $\downarrow$ 3111 2016 3707 5100 3779 2576 4731 8660 4542 3226 5921 13522
time $\downarrow$ 36’ 27’ 31’ 31’ 49’ 37’ 43’ 49’ 58’ 49’ 57’ 68’
Figure 13: PSNR/GFLOPs and PSNR/$||w||_{0}$ plots for aggregated results with TensoRF.
Table 7: Aggregated results on Instant-NGP
Dataset Metrics M=3 M=4 M=5
baseline Top-1 Top-2 Ens baseline Top-1 Top-2 Ens baseline Top-1 Top-2 Ens
[…]
SSIM $\uparrow$ 0.961 0.962 0.962 0.963 0.962 0.963 0.964 0.964 0.963 0.963 0.965 0.966
LPIPS $\downarrow$ 0.027 0.049 0.047 0.047 0.026 0.046 0.045 0.044 0.025 0.045 0.043 0.042
$\|w\|_{0}\downarrow$ 20 13 15 18 26 14 19 22 31 17 21 26
GFLOPs $\downarrow$ 42 66 111 122 44 69 117 165 46 71 123 208
time $\downarrow$ 7’ 23’ 26’ 27’ 9’ 27’ 30’ 30’ 11’ 32’ 34’ 37’
[…]
SSIM $\uparrow$ 0.981 0.981 0.982 0.982 0.982 0.983 0.983 0.984 0.983 0.983 0.984 0.985
LPIPS $\downarrow$ 0.012 0.024 0.024 0.023 0.011 0.023 0.023 0.022 0.010 0.023 0.022 0.020
$\|w\|_{0}\downarrow$ 20 15 16 18 26 16 19 22 30 17 20 27
GFLOPs $\downarrow$ 20 46 63 71 27 48 69 96 28 52 72 123
time $\downarrow$ 7’ 24’ 25’ 27’ 8’ 28’ 29’ 32’ 10’ 31’ 33’ 34’
[…]
SSIM $\uparrow$ 0.919 0.920 0.918 0.922 0.922 0.924 0.928 0.929 0.924 0.927 0.929 0.930
LPIPS $\downarrow$ 0.107 0.131 0.130 0.129 0.105 0.127 0.125 0.124 0.101 0.125 0.124 0.121
$\|w\|_{0}\downarrow$ 45 27 35 40 65 31 39 55 88 37 47 70
GFLOPs $\downarrow$ 210 220 498 590 215 231 522 785 211 229 531 1003
time $\downarrow$ 9’ 30’ 30’ 32’ 11’ 33’ 34’ 34’ 14’ 37’ 39’ 45’
[…]
SSIM $\uparrow$ 0.764 0.762 0.772 0.772 0.763 0.765 0.777 0.776 0.764 0.763 0.777 0.778
LPIPS $\downarrow$ 0.134 0.244 0.241 0.240 0.130 0.243 0.238 0.237 0.128 0.239 0.237 0.234
$\|w\|_{0}\downarrow$ 72 37 57 67 106 57 62 102 152 68 75 148
GFLOPs $\downarrow$ 539 761 1227 1607 529 882 1272 2172 573 887 1391 2712
time $\downarrow$ 14’ 20’ 22’ 22’ 18’ 28’ 30’ 29’ 24’ 25’ 46’ 42’
Figure 14: PSNR/GFLOPs and PSNR/$||w||_{0}$ plots for aggregated results with Instant-NGP.
Table 8: DVGO results on Synthetic-NeRF
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Lego 35.99 0.981 0.009 9 224 35.85 0.980 0.009 13 324 35.52 0.978 0.010 26 929
Mic 35.47 0.990 0.006 7 85 36.24 0.991 0.005 9 127 36.68 0.992 0.005 18 845
Chair 35.70 0.985 0.009 9 96 35.95 0.986 0.009 12 175 36.01 0.986 0.009 23 703
Hotdog 37.09 0.982 0.013 11 436 36.89 0.981 0.014 15 588 36.67 0.981 0.015 30 820
Ficus 33.90 0.983 0.012 8 121 33.80 0.982 0.013 12 145 33.86 0.982 0.014 22 773
Drums 25.81 0.935 0.051 8 258 25.97 0.937 0.049 11 320 25.94 0.936 0.049 20 1048
Materials 29.50 0.948 0.025 12 731 29.33 0.945 0.029 17 975 29.07 0.940 0.035 32 806
Ship 30.27 0.895 0.083 13 1107 30.56 0.898 0.074 18 1412 30.59 0.896 0.069 35 790
avg 32.97 0.962 0.026 10 382 33.07 0.963 0.025 13 508 33.04 0.961 0.026 26 839
Top-1 Lego 36.40 0.982 0.008 13 200 36.45 0.982 0.008 15 252 36.52 0.983 0.008 20 311
Mic 35.34 0.990 0.006 4 74 36.08 0.991 0.005 5 96 36.47 0.992 0.004 6 119
Chair 35.81 0.984 0.009 6 81 36.01 0.985 0.008 6 108 36.06 0.985 0.008 9 151
Hotdog 37.35 0.982 0.013 18 382 37.39 0.982 0.013 23 491 37.35 0.982 0.013 28 574
Ficus 34.37 0.983 0.012 13 115 34.47 0.984 0.012 15 136 34.45 0.984 0.012 20 147
Drums 25.80 0.932 0.054 10 251 25.91 0.934 0.051 11 291 26.00 0.935 0.049 16 327
Materials 29.95 0.952 0.025 39 643 29.91 0.952 0.025 49 824 29.72 0.950 0.026 59 963
Ship 30.32 0.892 0.082 29 994 30.53 0.895 0.076 40 1247 30.75 0.897 0.071 47 1395
avg 33.17 0.962 0.026 17 342 33.34 0.963 0.025 20 431 33.41 0.964 0.024 26 499
Top-2 Lego 36.59 0.983 0.008 20 378 36.80 0.983 0.008 26 476 36.89 0.983 0.007 31 587
Mic 35.49 0.990 0.005 6 140 36.43 0.992 0.004 9 181 36.98 0.993 0.004 8 225
Chair 36.16 0.986 0.008 8 153 36.41 0.987 0.007 11 204 36.54 0.987 0.007 12 285
Hotdog 37.50 0.983 0.013 27 721 37.68 0.984 0.011 38 926 37.78 0.984 0.011 48 1083
Ficus 34.20 0.983 0.012 22 217 34.28 0.983 0.011 21 256 34.45 0.984 0.011 26 277
Drums 25.98 0.936 0.050 15 475 26.05 0.937 0.048 22 548 26.21 0.938 0.046 21 616
Materials 29.98 0.953 0.023 53 1214 29.95 0.953 0.023 74 1555 29.99 0.954 0.022 87 1815
Ship 30.57 0.897 0.084 47 1878 30.78 0.900 0.076 60 2354 31.07 0.901 0.071 80 2631
avg 33.31 0.964 0.025 25 647 33.53 0.965 0.024 33 813 33.74 0.965 0.022 39 940
Ens Lego 36.54 0.983 0.008 27 534 36.75 0.983 0.007 47 895 36.82 0.984 0.007 76 1377
Mic 35.45 0.990 0.006 7 198 36.22 0.992 0.005 13 340 36.77 0.992 0.004 21 529
Chair 36.24 0.986 0.008 13 215 36.53 0.987 0.008 23 384 36.67 0.988 0.007 36 670
Hotdog 37.72 0.984 0.011 39 1019 37.79 0.985 0.011 75 1742 37.85 0.985 0.011 127 2542
Ficus 34.14 0.983 0.010 24 306 34.42 0.984 0.010 42 482 34.61 0.985 0.009 64 651
Drums 26.09 0.936 0.050 20 671 26.21 0.937 0.048 35 1031 26.39 0.940 0.045 56 1445
Materials 30.05 0.954 0.021 71 1715 30.11 0.955 0.019 136 2925 30.19 0.956 0.019 227 4262
Ship 30.53 0.897 0.081 53 2653 30.82 0.900 0.075 102 4428 31.06 0.903 0.070 169 6176
avg 33.35 0.964 0.025 32 914 33.61 0.965 0.023 59 1529 33.79 0.966 0.022 97 2206
Table 9: Per-scene results on NSVF with DVGO
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Bike 38.13 0.991 0.004 16 269 37.70 0.990 0.004 32 369 37.25 0.989 0.005 48 448
Lifestyle 34.27 0.969 0.021 20 369 34.34 0.969 0.021 42 505 34.24 0.968 0.022 65 619
Palace 36.59 0.974 0.015 21 189 37.01 0.976 0.013 43 264 36.90 0.975 0.012 66 319
Robot 37.24 0.993 0.004 31 385 37.07 0.992 0.004 64 585 36.56 0.991 0.004 100 730
Spaceship 36.53 0.983 0.016 58 399 35.63 0.979 0.022 113 536 34.78 0.974 0.028 179 662
Streamtrain 36.17 0.985 0.011 37 263 35.53 0.982 0.013 79 369 34.87 0.978 0.017 132 473
Toad 34.67 0.977 0.018 20 182 35.28 0.980 0.014 40 252 35.42 0.981 0.013 61 314
Wineholder 31.08 0.959 0.025 30 487 31.54 0.963 0.022 64 731 31.67 0.963 0.022 106 945
avg 35.59 0.979 0.014 29 318 35.51 0.979 0.014 60 452 35.21 0.977 0.015 95 564
Top-1 Bike 39.68 0.993 0.003 13 160 39.60 0.993 0.003 16 213 39.55 0.993 0.003 22 255
Lifestyle 35.25 0.971 0.018 15 307 35.27 0.972 0.016 21 416 35.37 0.973 0.015 25 467
Palace 37.53 0.977 0.010 21 426 37.97 0.980 0.010 26 549 38.24 0.981 0.008 28 681
Robot 38.79 0.995 0.003 13 158 39.07 0.995 0.002 16 205 39.11 0.995 0.002 23 245
Spaceship 38.43 0.988 0.009 20 238 38.43 0.988 0.010 24 296 38.17 0.987 0.011 33 356
Streamtrain 37.39 0.990 0.007 36 337 37.35 0.990 0.007 50 449 37.27 0.990 0.007 53 533
Toad 35.32 0.979 0.015 11 257 35.88 0.982 0.012 16 434 36.58 0.984 0.010 19 539
Wineholder 32.18 0.965 0.019 11 241 32.41 0.967 0.018 13 303 32.67 0.969 0.016 18 364
avg 36.82 0.982 0.010 18 266 37.00 0.983 0.010 23 358 37.12 0.984 0.009 27 430
Top-2 Bike 40.09 0.994 0.002 18 303 40.08 0.994 0.002 25 403 40.01 0.994 0.002 30 481
Lifestyle 35.51 0.974 0.015 17 580 35.65 0.975 0.015 26 785 35.67 0.975 0.014 33 881
Palace 37.94 0.980 0.011 30 806 38.50 0.982 0.009 45 1036 38.80 0.983 0.008 53 1283
Robot 39.29 0.995 0.002 19 300 39.43 0.995 0.002 26 387 39.46 0.995 0.002 35 462
Spaceship 38.98 0.989 0.007 31 449 39.00 0.990 0.007 35 560 39.03 0.990 0.007 48 671
Streamtrain 37.62 0.991 0.007 40 638 37.57 0.991 0.007 66 848 37.54 0.991 0.007 95 1005
Toad 36.19 0.982 0.012 12 486 36.80 0.985 0.010 17 819 37.16 0.985 0.009 25 1016
Wineholder 32.43 0.967 0.018 16 456 32.83 0.970 0.015 18 572 33.06 0.971 0.014 27 686
avg 37.26 0.984 0.009 23 502 37.48 0.985 0.008 32 676 37.59 0.986 0.008 43 811
Ens Bike 39.84 0.993 0.002 24 428 40.00 0.994 0.002 44 758 39.95 0.994 0.002 70 1130
Lifestyle 35.50 0.974 0.015 23 819 35.67 0.976 0.013 43 1477 35.75 0.976 0.012 68 2069
Palace 37.98 0.980 0.010 38 1138 38.66 0.983 0.008 69 1948 39.03 0.984 0.007 115 3012
Robot 39.08 0.995 0.002 24 423 39.33 0.995 0.002 45 727 39.43 0.995 0.002 74 1084
Spaceship 39.03 0.990 0.007 44 635 39.15 0.990 0.007 81 1052 39.22 0.990 0.007 137 1576
Streamtrain 37.49 0.991 0.007 60 901 37.52 0.991 0.007 115 1595 37.50 0.990 0.007 192 2359
Toad 36.28 0.983 0.011 24 686 37.05 0.986 0.009 50 1540 37.53 0.987 0.008 86 2385
Wineholder 32.22 0.967 0.018 19 644 32.70 0.970 0.016 35 1075 33.06 0.972 0.014 56 1610
avg 37.18 0.984 0.009 32 709 37.51 0.985 0.008 60 1271 37.68 0.986 0.007 100 1903
Table 10: Per-scene results on TanksAndTemple with DVGO
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Ignatius 28.39 0.95 0.076 20 1588 28.55 0.951 0.07 35 1755 27.86 0.873 0.207 84 3490
Caterpillar 26.06 0.911 0.137 9 1298 26.08 0.912 0.129 31 1012 27.68 0.924 0.102 59 1592
Family 33.76 0.964 0.057 23 1293 33.75 0.964 0.053 20 1647 26.13 0.914 0.119 51 2561
Barn 27.28 0.853 0.258 16 822 27.61 0.860 0.236 20 1588 28.52 0.951 0.068 89 3088
Truck 27.45 0.915 0.126 18 1493 27.55 0.920 0.115 47 1735 34.43 0.97 0.039 86 2599
avg 28.59 0.919 0.131 17 1299 28.71 0.921 0.121 31 1547 28.92 0.926 0.107 74 2666
Top-1 Ignatius 28.18 0.947 0.073 11 796 28.28 0.948 0.071 12 926 28.36 0.949 0.070 14 1044
Caterpillar 26.45 0.914 0.121 11 1265 26.55 0.914 0.119 14 1462 26.59 0.916 0.117 14 1716
Family 34.5 0.968 0.049 11 1140 34.67 0.969 0.046 15 1457 34.75 0.97 0.044 18 1714
Barn 27.85 0.859 0.223 12 1384 27.98 0.865 0.21 16 1681 28.18 0.868 0.204 18 1656
Truck 27.69 0.918 0.11 12 1527 27.78 0.921 0.105 15 1789 27.8 0.921 0.104 18 2002
avg 28.93 0.921 0.115 11 1223 29.05 0.923 0.110 15 1463 29.14 0.925 0.108 16 1626
Top-2 Ignatius 28.21 0.949 0.074 17 1505 28.27 0.950 0.071 21 1747 28.27 0.951 0.069 28 1969
Caterpillar 26.66 0.918 0.12 16 2390 26.72 0.920 0.117 19 2759 26.78 0.921 0.114 24 3235
Family 34.75 0.97 0.048 17 2155 34.92 0.972 0.045 20 2750 35 0.972 0.043 24 3232
Barn 28.14 0.866 0.212 19 2616 28.27 0.871 0.202 26 3173 28.42 0.874 0.195 27 3122
Truck 27.75 0.921 0.11 19 2885 27.84 0.924 0.105 23 3377 27.88 0.925 0.101 26 3773
avg 29.10 0.925 0.113 18 2310 29.21 0.927 0.108 22 2761 29.27 0.929 0.105 26 3066
Ens Ignatius 28.26 0.95 0.076 21 2126 28.36 0.953 0.072 37 3286 28.4 0.954 0.07 57 4622
Caterpillar 26.66 0.919 0.122 20 3377 26.85 0.923 0.118 30 5189 26.94 0.925 0.113 46 7593
Family 34.72 0.971 0.047 28 3045 34.93 0.973 0.044 50 5173 35.06 0.974 0.042 82 7588
Barn 28.13 0.868 0.209 27 3695 28.32 0.874 0.197 48 5968 28.51 0.879 0.19 68 7328
Truck 27.76 0.923 0.111 27 4075 27.89 0.927 0.104 46 6350 27.95 0.93 0.101 71 8857
avg 29.11 0.926 0.113 25 3263 29.27 0.930 0.107 43 5193 29.37 0.932 0.103 65 7198
Table 11: Per-scene results on LLFF with DVGO
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Fern 24.89 0.816 0.166 28 1062 24.76 0.816 0.161 49 1382 24.58 0.810 0.160 95 1859
Flower 27.51 0.850 0.132 16 1018 27.65 0.854 0.127 28 1295 27.67 0.856 0.122 53 1690
Fortress 29.91 0.859 0.121 8 640 29.42 0.831 0.143 16 907 29.38 0.821 0.149 36 1407
Horns 27.55 0.863 0.142 19 1017 27.83 0.874 0.129 32 1258 27.86 0.878 0.120 60 1641
Leaves 20.96 0.737 0.181 21 915 21.06 0.744 0.167 36 1187 21.08 0.748 0.156 67 1547
Orchids 20.40 0.683 0.186 29 1354 20.51 0.683 0.181 49 1684 20.42 0.680 0.179 86 2096
Room 31.36 0.944 0.107 13 922 31.75 0.947 0.101 22 1154 31.65 0.937 0.114 43 1613
Trex 27.04 0.904 0.094 16 877 27.34 0.910 0.089 29 1136 27.24 0.913 0.085 57 1579
avg 26.20 0.832 0.141 19 976 26.29 0.832 0.137 33 1251 26.24 0.831 0.136 62 1679
Top-1 Fern 24.31 0.798 0.162 16 932 24.45 0.804 0.149 27 1299 24.48 0.806 0.150 39 1518
Flower 27.44 0.846 0.115 12 872 27.53 0.851 0.107 19 1175 27.72 0.858 0.099 26 1518
Fortress 29.71 0.836 0.131 7 618 29.96 0.847 0.113 10 812 29.93 0.845 0.116 16 1531
Horns 27.68 0.861 0.129 12 957 28.08 0.877 0.107 18 1277 28.29 0.883 0.101 24 1460
Leaves 20.76 0.735 0.152 14 814 20.91 0.745 0.139 22 1109 20.95 0.747 0.131 31 1615
Orchids 19.68 0.658 0.165 17 986 19.71 0.659 0.159 29 1423 19.73 0.655 0.158 37 1660
Room 31.64 0.943 0.098 9 840 32.08 0.947 0.092 13 1193 32.40 0.947 0.093 17 1522
Trex 27.17 0.903 0.085 10 734 27.71 0.914 0.073 15 1006 27.92 0.918 0.067 21 1291
avg 26.05 0.823 0.130 12 844 26.30 0.830 0.117 19 1162 26.43 0.832 0.115 26 1514
Top-2 Fern 24.34 0.801 0.156 30 1555 24.57 0.811 0.148 43 2159 24.36 0.808 0.144 62 2815
Flower 27.83 0.858 0.108 21 1454 27.86 0.861 0.105 27 1953 27.97 0.864 0.100 37 2467
Fortress 30.34 0.867 0.098 14 1030 30.18 0.859 0.105 16 1349 30.18 0.860 0.105 28 2123
Horns 28.05 0.874 0.118 24 1596 28.42 0.883 0.108 27 2123 28.69 0.890 0.100 30 2609
Leaves 20.74 0.737 0.151 27 1358 20.97 0.750 0.141 38 1844 20.99 0.752 0.133 51 2351
Orchids 19.64 0.654 0.166 31 1644 19.66 0.656 0.159 35 2366 19.70 0.658 0.156 55 3017
Room 32.25 0.950 0.087 17 1401 32.51 0.952 0.087 22 1983 32.96 0.955 0.081 29 2506
Trex 27.45 0.910 0.079 18 1224 27.90 0.919 0.072 22 1671 28.07 0.923 0.068 32 2174
avg 26.33 0.831 0.120 23 1408 26.51 0.837 0.116 29 1931 26.62 0.839 0.111 40 2508
Ens Fern 24.55 0.806 0.155 38 1870 24.69 0.811 0.144 83 3444 24.72 0.813 0.143 162 5582
Flower 27.70 0.858 0.113 26 1748 27.85 0.863 0.107 53 3115 27.95 0.864 0.102 97 4892
Fortress 30.22 0.869 0.094 16 1239 30.43 0.874 0.081 32 2151 30.24 0.874 0.096 72 4208
Horns 27.94 0.876 0.117 31 1919 28.29 0.890 0.099 64 3385 28.56 0.891 0.087 115 5174
Leaves 20.81 0.739 0.160 31 1633 20.95 0.749 0.146 65 2940 21.01 0.750 0.137 123 4661
Orchids 19.86 0.666 0.173 39 1977 19.90 0.667 0.167 84 3774 19.95 0.667 0.163 154 5983
Room 32.15 0.951 0.087 21 1685 32.57 0.955 0.080 45 3162 33.03 0.965 0.075 83 4969
Trex 27.20 0.908 0.082 23 1471 27.56 0.917 0.074 49 2664 27.72 0.917 0.070 94 4310
avg 26.31 0.834 0.123 28 1693 26.53 0.841 0.112 59 3079 26.65 0.843 0.109 113 4972
Table 12: Per-scene results on Synthetic-NeRF with TensorF
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Lego 36.85 0.986 0.006 15 376 36.73 0.986 0.006 25 474 36.70 0.986 0.007 38 579
Mic 34.98 0.994 0.005 8 188 34.80 0.994 0.005 13 259 34.45 0.994 0.005 19 307
Chair 30.89 0.879 0.088 13 154 30.80 0.879 0.078 22 198 30.70 0.871 0.079 33 255
Hotdog 26.01 0.942 0.050 22 684 25.91 0.941 0.050 37 818 25.44 0.926 0.059 61 982
Ficus 37.51 0.981 0.014 17 348 37.42 0.981 0.013 29 406 37.28 0.980 0.014 44 455
Drums 35.98 0.995 0.003 15 371 35.97 0.995 0.003 25 454 35.83 0.995 0.003 31 514
Materials 30.06 0.937 0.039 17 1359 29.86 0.933 0.044 29 1690 29.62 0.930 0.050 46 2007
Ship 34.13 0.986 0.011 17 1318 34.02 0.985 0.013 30 1723 33.84 0.984 0.015 46 1988
avg 33.30 0.962 0.027 15 600 33.19 0.962 0.027 26 753 32.98 0.958 0.029 40 886
Top-1 Lego 36.70 0.983 0.007 10 330 36.82 0.986 0.007 17 395 37.02 0.984 0.006 22 460
Mic 34.93 0.989 0.008 6 157 35.13 0.989 0.007 9 196 35.56 0.990 0.006 13 224
Chair 36.00 0.985 0.009 9 122 36.24 0.985 0.009 12 157 36.34 0.986 0.009 19 192
Hotdog 38.00 0.984 0.011 14 605 38.05 0.984 0.012 22 696 37.95 0.983 0.012 30 832
Ficus 34.31 0.983 0.012 12 323 34.43 0.986 0.012 20 379 34.57 0.984 0.011 30 412
Drums 25.99 0.936 0.050 9 292 26.28 0.940 0.046 15 386 26.17 0.938 0.048 23 437
Materials 30.52 0.957 0.021 12 1166 30.58 0.956 0.022 20 1428 30.53 0.956 0.022 26 1681
Ship 31.14 0.902 0.082 12 1167 31.14 0.902 0.078 15 1492 31.32 0.901 0.076 30 1615
avg 33.45 0.965 0.025 11 520 33.58 0.966 0.024 16 641 33.68 0.965 0.024 24 732
Top-2 Lego 37.38 0.985 0.006 14 605 37.26 0.985 0.006 23 723 37.59 0.986 0.006 31 840
Mic 35.03 0.989 0.007 9 285 35.33 0.990 0.006 12 355 35.48 0.991 0.006 19 404
Chair 36.55 0.988 0.007 13 220 36.61 0.988 0.007 18 282 36.67 0.988 0.008 26 345
Hotdog 38.01 0.985 0.011 20 1115 38.13 0.985 0.010 30 1281 38.29 0.986 0.010 44 1530
Ficus 35.42 0.987 0.008 17 592 35.65 0.987 0.008 24 693 35.60 0.985 0.008 38 753
Drums 26.08 0.937 0.050 12 534 26.26 0.941 0.045 21 706 26.33 0.941 0.045 31 799
Materials 30.99 0.961 0.018 16 2155 31.17 0.963 0.017 24 2637 31.04 0.962 0.018 37 3103
Ship 31.21 0.903 0.081 17 2155 31.54 0.906 0.075 26 2754 31.75 0.907 0.072 37 2979
avg 33.83 0.967 0.024 15 958 33.99 0.968 0.022 23 1179 34.09 0.968 0.021 33 1344
Ens Lego 37.27 0.985 0.006 18 839 37.35 0.986 0.006 31 1120 37.64 0.986 0.006 49 1402
Mic 35.10 0.989 0.007 11 397 35.25 0.989 0.007 19 531 35.58 0.990 0.006 30 666
Chair 36.52 0.987 0.008 16 307 36.62 0.988 0.008 27 411 36.75 0.988 0.007 43 516
Hotdog 38.02 0.985 0.011 23 1545 38.15 0.985 0.011 41 2061 38.22 0.985 0.010 67 2578
Ficus 34.40 0.984 0.010 20 822 34.67 0.985 0.009 35 1097 34.95 0.986 0.009 55 1374
Drums 26.16 0.937 0.051 14 741 26.27 0.937 0.050 25 989 26.34 0.938 0.051 39 1239
Materials 30.64 0.958 0.021 19 2981 30.75 0.959 0.020 33 3975 30.78 0.959 0.020 53 4971
Ship 31.15 0.902 0.079 20 2981 31.54 0.906 0.074 35 3976 31.78 0.908 0.070 56 4971
avg 33.66 0.966 0.024 18 1327 33.83 0.967 0.023 31 1770 34.00 0.968 0.022 49 2214
Table 13: Per-scene results on NSVF with TensoRF
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Bike 39.48 0.996 0.003 14 304 39.25 0.996 0.003 23 362 38.59 0.995 0.004 35 421
Lifestyle 34.76 0.975 0.015 18 680 34.71 0.974 0.016 30 833 34.53 0.973 0.019 45 990
Palace 38.20 0.983 0.008 15 541 38.51 0.984 0.008 25 671 38.60 0.984 0.007 38 803
Robot 38.86 0.997 0.002 13 229 38.98 0.997 0.002 21 286 38.81 0.997 0.002 30 347
Spaceship 38.68 0.993 0.006 15 487 38.23 0.990 0.010 25 593 37.51 0.986 0.017 39 689
Streamtrain 38.03 0.993 0.008 20 536 37.99 0.994 0.007 33 673 37.61 0.993 0.006 51 806
Toad 35.43 0.950 0.029 21 564 35.79 0.947 0.029 36 710 35.81 0.949 0.028 55 872
Wineholder 31.84 0.968 0.021 17 474 32.11 0.969 0.019 29 601 32.11 0.968 0.020 44 721
avg 36.91 0.982 0.011 17 477 36.95 0.981 0.012 28 591 36.70 0.981 0.013 42 706
Top-1 Bike 39.97 0.993 0.002 9 264 40.06 0.993 0.002 16 321 40.16 0.994 0.002 24 362
Lifestyle 35.23 0.971 0.018 11 598 35.09 0.970 0.019 20 719 35.23 0.972 0.017 29 823
Palace 37.23 0.978 0.012 9 469 38.00 0.980 0.011 18 584 38.64 0.982 0.009 25 651
Robot 38.61 0.994 0.003 9 199 38.89 0.994 0.003 16 238 39.04 0.995 0.003 24 271
Spaceship 39.34 0.990 0.007 9 415 39.50 0.990 0.007 18 514 39.56 0.990 0.007 28 588
Streamtrain 38.49 0.992 0.005 12 469 38.57 0.992 0.005 22 565 38.59 0.992 0.005 30 642
Toad 35.01 0.978 0.015 12 463 34.83 0.977 0.017 19 585 35.51 0.981 0.014 34 682
Wineholder 32.00 0.965 0.020 9 437 32.53 0.967 0.019 20 505 32.46 0.968 0.018 29 577
avg 36.98 0.983 0.010 10 414 37.18 0.983 0.010 19 504 37.40 0.984 0.009 28 575
Top-2 Bike 40.44 0.994 0.002 13 482 40.61 0.994 0.002 23 586 40.49 0.994 0.002 34 660
Lifestyle 35.52 0.973 0.017 15 1101 35.70 0.974 0.016 25 1322 35.75 0.974 0.016 39 1513
Palace 38.34 0.982 0.009 15 862 39.02 0.984 0.008 24 1071 39.35 0.986 0.007 34 1194
Robot 39.37 0.995 0.002 15 361 39.66 0.995 0.002 23 432 39.88 0.996 0.002 34 492
Spaceship 39.90 0.991 0.006 13 764 39.91 0.991 0.006 23 944 39.87 0.991 0.006 35 1079
Streamtrain 38.65 0.992 0.005 18 862 38.48 0.992 0.005 32 1038 38.75 0.993 0.005 40 1177
Toad 35.95 0.982 0.012 18 850 36.30 0.984 0.011 31 1074 36.60 0.985 0.010 48 1251
Wineholder 32.48 0.968 0.019 16 803 32.81 0.970 0.017 26 927 33.12 0.972 0.016 38 1056
avg 37.58 0.985 0.009 15 761 37.81 0.986 0.008 26 924 37.98 0.986 0.008 38 1053
Ens Bike 40.41 0.994 0.002 18 671 40.66 0.994 0.002 30 897 40.76 0.994 0.002 46 1124
Lifestyle 35.55 0.973 0.017 20 1525 35.79 0.975 0.015 34 2035 35.95 0.976 0.014 55 2545
Palace 38.40 0.983 0.009 19 1194 38.99 0.985 0.008 33 1593 39.41 0.986 0.007 51 1992
Robot 39.35 0.995 0.002 19 503 39.66 0.995 0.002 33 672 39.83 0.996 0.002 52 843
Spaceship 39.85 0.991 0.006 17 1060 40.08 0.991 0.006 29 1415 40.18 0.992 0.005 46 1771
Streamtrain 38.38 0.992 0.006 23 1197 38.48 0.992 0.005 40 1598 38.58 0.992 0.005 64 2000
Toad 36.07 0.983 0.012 23 1179 36.55 0.985 0.010 40 1573 36.89 0.986 0.009 64 1968
Wineholder 32.28 0.967 0.020 21 1114 32.57 0.969 0.019 37 1486 33.06 0.972 0.016 58 1860
avg 37.54 0.985 0.009 20 1055 37.85 0.986 0.008 35 1408 38.08 0.987 0.008 54 1763
Table 14: Per-scene results on TanksAndTemple with TensoRF
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Caterpillar 26.16 0.842 0.191 3 3155 25.83 0.847 0.177 5 3674 25.82 0.845 0.172 7 4557
Family 34.20 0.963 0.059 5 3071 34.21 0.963 0.053 8 3578 33.94 0.963 0.050 13 4366
Barn 27.59 0.949 0.076 0 1841 27.68 0.951 0.066 1 2489 27.68 0.951 0.064 1 2917
Truck 26.84 0.859 0.223 3 2386 26.60 0.852 0.227 4 3004 26.47 0.845 0.214 6 3743
Ignatius 28.47 0.915 0.132 3 1441 28.42 0.917 0.126 6 1830 28.29 0.918 0.116 10 2253
avg 28.65 0.906 0.136 3 2379 28.55 0.906 0.130 5 2915 28.44 0.905 0.123 7 3567
Top-1 Caterpillar 26.46 0.914 0.125 5 2018 26.45 0.914 0.122 7 2423 26.41 0.916 0.116 9 2839
Family 34.08 0.965 0.050 3 1548 34.08 0.966 0.048 4 1775 34.27 0.968 0.044 5 2318
Barn 27.75 0.868 0.194 6 3183 27.75 0.874 0.180 8 3377 27.57 0.872 0.187 11 3455
Truck 27.28 0.915 0.119 5 2703 27.15 0.917 0.110 8 3317 27.26 0.916 0.113 11 3621
Ignatius 28.36 0.946 0.074 5 1303 28.44 0.947 0.069 6 1481 28.40 0.946 0.069 9 1723
avg 28.79 0.922 0.112 5 2151 28.77 0.924 0.106 7 2475 28.78 0.924 0.106 9 2791
Top-2 Caterpillar 26.65 0.917 0.123 8 3712 26.65 0.918 0.119 11 4452 26.67 0.920 0.112 16 5211
Family 34.28 0.967 0.051 5 2845 34.33 0.968 0.049 7 3258 34.58 0.970 0.043 10 4256
Barn 28.06 0.873 0.190 10 5866 28.22 0.880 0.173 14 6215 28.31 0.883 0.167 19 6347
Truck 27.43 0.917 0.118 8 4979 27.48 0.921 0.110 13 6104 27.56 0.922 0.105 18 6657
Ignatius 28.57 0.948 0.073 7 2394 28.58 0.949 0.070 10 2716 28.60 0.950 0.067 14 3157
avg 29.00 0.924 0.111 8 3959 29.05 0.927 0.104 11 4549 29.14 0.929 0.099 15 5126
Ens Caterpillar 26.64 0.917 0.127 11 5138 26.75 0.919 0.127 48 6853 26.77 0.919 0.126 79 8571
Family 34.11 0.966 0.053 8 3942 34.16 0.967 0.055 45 5260 34.32 0.968 0.052 74 6581
Barn 28.06 0.874 0.189 13 8118 28.21 0.878 0.183 54 10827 28.32 0.882 0.176 89 13541
Truck 27.35 0.917 0.122 11 6889 27.40 0.919 0.122 51 9188 27.45 0.921 0.119 84 11491
Ignatius 28.59 0.949 0.075 10 3320 28.67 0.951 0.077 45 4430 28.70 0.952 0.074 74 5545
avg 28.95 0.925 0.113 11 5481 29.04 0.927 0.113 49 7311 29.11 0.928 0.109 80 9146
Table 15: Per-scene results on LLFF with TensoRF
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Fern 24.98 0.801 0.165 9 4124 25.01 0.798 0.162 13 4766 25.04 0.801 0.152 18 5671
Flower 28.41 0.863 0.108 9 3413 28.38 0.865 0.099 13 4048 28.33 0.863 0.097 19 5054
Fortress 31.18 0.892 0.079 7 3123 31.27 0.895 0.072 11 3922 31.43 0.899 0.066 15 4872
Horns 28.04 0.871 0.125 9 1980 28.30 0.880 0.110 12 2523 28.41 0.886 0.100 18 2940
Leaves 21.16 0.742 0.165 8 3152 21.07 0.742 0.150 12 3690 20.95 0.738 0.144 17 4521
Orchids 19.84 0.647 0.200 12 3658 19.81 0.645 0.193 17 4472 19.75 0.636 0.199 24 5241
Room 31.78 0.948 0.089 10 2597 32.22 0.951 0.079 15 3357 32.23 0.951 0.076 21 3861
Trex 27.43 0.906 0.087 9 2844 27.70 0.911 0.079 13 3457 27.55 0.910 0.078 18 4177
avg 26.60 0.834 0.127 9 3111 26.72 0.836 0.118 13 3779 26.71 0.835 0.114 19 4542
Top-1 Fern 24.76 0.792 0.166 5 2703 24.78 0.800 0.166 8 3438 24.76 0.800 0.167 12 4223
Flower 28.57 0.868 0.089 5 2155 28.60 0.873 0.089 8 2825 28.78 0.875 0.079 11 3500
Fortress 31.13 0.893 0.075 4 1928 31.17 0.898 0.071 8 2517 31.23 0.896 0.070 8 3526
Horns 28.09 0.873 0.113 5 1251 28.44 0.882 0.100 6 1676 28.42 0.872 0.101 10 1956
Leaves 21.30 0.748 0.151 5 2033 21.17 0.746 0.142 7 2565 21.05 0.740 0.139 13 3010
Orchids 19.89 0.656 0.168 6 2449 19.80 0.648 0.173 9 3129 19.76 0.645 0.175 15 3770
Room 31.97 0.948 0.083 5 1578 32.34 0.951 0.076 10 2218 32.26 0.950 0.079 11 2713
Trex 26.94 0.900 0.091 4 2031 27.56 0.910 0.082 7 2237 27.55 0.912 0.082 10 3108
avg 26.58 0.835 0.117 5 2016 26.73 0.839 0.112 6 2576 26.73 0.836 0.111 11 3226
Top-2 Fern 24.88 0.797 0.156 6 4974 24.87 0.795 0.157 10 6320 24.89 0.796 0.156 16 7754
Flower 28.84 0.873 0.086 6 3965 28.87 0.875 0.082 10 5192 29.05 0.878 0.078 15 6425
Fortress 31.43 0.899 0.067 6 3545 31.50 0.900 0.063 10 4622 31.68 0.902 0.056 14 6473
Horns 28.55 0.884 0.103 6 2297 28.82 0.891 0.093 10 3074 29.08 0.899 0.085 15 3584
Leaves 21.37 0.754 0.145 6 3740 21.20 0.752 0.138 10 4713 21.25 0.755 0.135 16 5525
Orchids 19.94 0.661 0.164 7 4505 19.87 0.657 0.162 13 5751 19.90 0.658 0.161 20 6922
Room 32.46 0.952 0.077 7 2899 32.97 0.955 0.070 11 4073 33.06 0.955 0.068 17 4978
Trex 27.44 0.909 0.080 6 3734 27.76 0.915 0.073 9 4106 27.79 0.914 0.072 12 5704
avg 26.86 0.841 0.110 6 3707 26.98 0.843 0.105 10 4731 27.09 0.845 0.101 16 5921
Ens Fern 24.80 0.790 0.170 8 6841 24.83 0.789 0.165 13 11566 24.80 0.788 0.160 22 17709
Flower 28.79 0.873 0.091 8 5454 28.88 0.875 0.085 13 9503 29.13 0.880 0.080 21 14676
Fortress 31.38 0.899 0.069 8 4876 31.50 0.902 0.062 14 8458 31.65 0.904 0.056 21 14783
Horns 28.50 0.885 0.105 8 3161 28.88 0.895 0.090 14 5627 29.10 0.904 0.080 22 8186
Leaves 21.34 0.751 0.154 8 5146 21.26 0.753 0.142 13 8629 21.26 0.757 0.135 22 12623
Orchids 20.04 0.660 0.171 9 6196 20.05 0.659 0.167 17 10525 20.04 0.660 0.164 27 15807
Room 32.38 0.952 0.079 9 3988 32.91 0.956 0.070 16 7455 33.08 0.957 0.066 25 11368
Trex 27.15 0.906 0.088 8 5138 27.70 0.915 0.075 14 7516 27.78 0.916 0.071 22 13028
avg 26.80 0.840 0.116 8 5100 27.00 0.843 0.107 14 8660 27.10 0.846 0.101 23 13522
Table 16: Per-scene results on Synthetic-NeRF with Instant-NGP
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Lego 35.88 0.980 0.009 12 23 35.98 0.981 0.008 14 29 36.09 0.981 0.008 15 30
Mic 36.25 0.990 0.008 10 18 36.27 0.991 0.007 12 20 36.66 0.991 0.007 13 22
Ship 30.69 0.892 0.080 33 130 30.58 0.893 0.075 46 137 30.59 0.894 0.073 51 145
Chair 35.57 0.982 0.010 5 13 35.76 0.980 0.009 6 13 35.80 0.983 0.009 6 13
Hotdog 37.35 0.981 0.015 7 40 37.48 0.981 0.014 8 40 37.60 0.982 0.014 8 42
Materials 29.64 0.947 0.032 45 59 29.76 0.949 0.030 63 59 29.84 0.951 0.028 82 60
Ficus 34.19 0.983 0.013 13 18 34.26 0.983 0.013 17 19 34.26 0.983 0.013 19 22
Drums 25.92 0.934 0.052 31 32 25.94 0.934 0.050 43 33 25.95 0.936 0.050 52 34
avg 33.19 0.961 0.027 20 42 33.25 0.962 0.026 26 44 33.35 0.963 0.025 31 46
Top-1 Lego 36.19 0.983 0.019 11 39 36.38 0.983 0.019 12 42 36.58 0.985 0.018 15 45
Mic 36.27 0.990 0.013 9 38 36.52 0.991 0.014 10 37 36.72 0.993 0.013 11 38
Ship 30.79 0.892 0.122 17 178 30.86 0.893 0.126 22 190 30.85 0.893 0.126 27 198
Chair 35.77 0.984 0.019 4 27 35.95 0.980 0.019 4 27 35.98 0.980 0.019 6 27
Hotdog 37.45 0.981 0.030 5 59 37.49 0.981 0.032 6 59 37.69 0.981 0.032 9 57
Materials 29.70 0.947 0.060 30 89 29.76 0.952 0.062 32 97 29.86 0.952 0.064 39 100
Ficus 34.29 0.983 0.019 11 45 34.46 0.985 0.019 10 45 34.76 0.985 0.019 10 45
Drums 25.98 0.934 0.074 21 58 25.99 0.935 0.072 19 58 25.99 0.935 0.071 22 57
avg 33.31 0.962 0.04 13 66 33.43 0.963 0.046 14 69 33.56 0.963 0.045 17 71
Top-2 Lego 36.32 0.981 0.020 12 66 36.37 0.982 0.019 14 66 36.47 0.983 0.018 15 73
Mic 36.70 0.991 0.014 9 64 36.85 0.991 0.014 11 64 37.00 0.992 0.013 12 65
Ship 30.38 0.892 0.129 20 294 30.56 0.894 0.126 28 312 30.79 0.897 0.122 31 342
Chair 35.87 0.982 0.021 4 45 35.91 0.983 0.019 5 45 36.05 0.985 0.018 6 50
Hotdog 37.80 0.982 0.032 5 97 37.82 0.982 0.032 6 102 38.02 0.983 0.029 7 98
Materials 29.97 0.948 0.066 36 151 30.18 0.952 0.061 47 167 30.36 0.954 0.058 53 165
Ficus 35.68 0.987 0.019 12 76 35.81 0.987 0.018 13 76 35.65 0.987 0.018 14 75
Drums 26.16 0.934 0.074 24 93 26.28 0.936 0.072 30 100 26.29 0.937 0.071 33 116
avg 33.61 0.962 0.047 15 111 33.72 0.964 0.045 19 117 33.83 0.965 0.043 21 123
Ens Lego 36.25 0.981 0.020 13 67 36.55 0.982 0.018 15 91 36.67 0.984 0.017 19 118
Mic 36.95 0.992 0.013 11 53 37.32 0.992 0.012 13 71 37.62 0.993 0.010 14 89
Ship 30.66 0.893 0.129 25 385 30.97 0.897 0.124 26 526 31.19 0.899 0.121 31 657
Chair 26.24 0.936 0.073 5 93 26.42 0.938 0.070 6 125 26.53 0.941 0.067 6 159
Hotdog 34.77 0.985 0.022 6 54 35.07 0.986 0.020 6 71 35.26 0.986 0.019 6 92
Materials 29.85 0.948 0.066 43 174 30.15 0.952 0.060 53 232 30.36 0.955 0.055 65 289
Ficus 37.74 0.982 0.031 14 112 37.94 0.984 0.029 16 153 38.11 0.984 0.027 18 197
Drums 35.77 0.985 0.021 28 36 36.13 0.984 0.018 35 49 36.33 0.987 0.017 44 63
avg 33.53 0.963 0.047 18 122 33.82 0.964 0.044 21 165 34.01 0.966 0.042 26 208
Table 17: Per-scene results on NSVF with Instant-NGP
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Bike 37.86 0.990 0.004 13 12 38.04 0.991 0.004 17 17 38.45 0.991 0.003 20 17
Lifestyle 34.80 0.968 0.019 24 31 34.95 0.970 0.018 30 49 35.05 0.970 0.017 35 51
Palace 37.66 0.978 0.010 27 31 38.18 0.981 0.008 31 35 38.38 0.982 0.008 32 37
Robot 37.02 0.994 0.004 20 13 37.44 0.994 0.004 26 14 37.50 0.994 0.004 32 17
Spaceship 36.74 0.984 0.013 12 17 36.77 0.984 0.013 13 21 36.76 0.985 0.013 16 23
Streamtrain 36.54 0.988 0.009 21 18 36.50 0.988 0.009 30 22 36.71 0.989 0.008 37 23
Toad 35.94 0.981 0.012 13 17 36.32 0.983 0.011 15 23 36.59 0.984 0.010 15 25
Wineholder 31.89 0.964 0.021 31 24 31.97 0.966 0.021 41 33 32.07 0.967 0.020 54 35
avg 36.06 0.981 0.012 20 20 36.27 0.982 0.011 26 27 36.44 0.983 0.010 30 28
Top-1 Bike 38.41 0.989 0.013 11 32 38.50 0.990 0.013 11 32 38.79 0.991 0.011 13 37
Lifestyle 35.06 0.968 0.043 15 81 35.25 0.970 0.040 15 86 35.29 0.971 0.039 17 88
Palace 37.42 0.978 0.019 19 62 37.46 0.981 0.017 21 67 37.66 0.981 0.017 22 68
Robot 37.60 0.995 0.009 17 32 37.72 0.995 0.009 18 32 37.75 0.995 0.009 21 36
Spaceship 37.22 0.985 0.023 9 39 37.33 0.985 0.024 11 40 37.48 0.985 0.024 12 41
Streamtrain 36.88 0.988 0.014 18 38 37.02 0.989 0.017 18 37 37.29 0.989 0.017 20 47
Toad 35.99 0.980 0.021 10 34 36.15 0.981 0.021 11 35 36.15 0.982 0.021 13 39
Wineholder 32.04 0.965 0.045 18 51 32.13 0.968 0.043 20 52 32.35 0.969 0.043 21 59
avg 36.33 0.981 0.024 15 46 36.44 0.982 0.023 16 48 36.59 0.983 0.023 17 52
Top-2 Bike 38.63 0.991 0.013 12 52 38.76 0.992 0.013 11 55 38.76 0.992 0.013 11 57
Lifestyle 35.36 0.968 0.041 16 88 35.49 0.970 0.039 18 98 35.66 0.971 0.037 21 103
Palace 37.74 0.978 0.019 20 87 38.18 0.980 0.017 23 94 38.56 0.982 0.016 24 101
Robot 37.86 0.996 0.009 18 55 38.01 0.996 0.009 19 57 38.15 0.996 0.009 22 59
Spaceship 37.58 0.985 0.024 6 47 37.83 0.986 0.024 10 60 38.23 0.987 0.022 11 64
Streamtrain 37.05 0.989 0.018 17 54 37.19 0.990 0.017 23 60 37.24 0.990 0.017 24 64
Toad 36.13 0.982 0.022 9 48 36.45 0.983 0.020 12 49 36.92 0.984 0.018 12 53
Wineholder 32.56 0.967 0.044 26 71 32.75 0.968 0.042 27 78 32.78 0.969 0.040 30 82
avg 36.61 0.982 0.024 16 63 36.83 0.983 0.023 18 69 37.04 0.984 0.022 20 73
Ens Bike 38.63 0.991 0.013 12 40 38.98 0.992 0.012 15 55 39.28 0.993 0.011 17 68
Lifestyle 35.40 0.968 0.040 20 148 35.71 0.971 0.037 22 197 35.90 0.971 0.035 27 253
Palace 37.94 0.979 0.018 20 100 38.45 0.982 0.016 24 136 38.87 0.984 0.015 27 173
Robot 37.72 0.994 0.009 19 40 38.16 0.996 0.009 25 54 38.46 0.995 0.009 29 69
Spaceship 37.58 0.986 0.024 10 58 38.03 0.987 0.022 12 78 38.40 0.988 0.021 13 98
Streamtrain 36.87 0.989 0.018 21 50 37.15 0.990 0.017 26 67 37.32 0.990 0.016 34 86
Toad 36.05 0.982 0.022 12 49 36.63 0.983 0.019 12 67 37.14 0.985 0.018 15 85
Wineholder 32.49 0.967 0.043 29 86 32.84 0.970 0.040 37 116 33.10 0.972 0.038 48 149
avg 36.58 0.982 0.023 18 71 36.99 0.984 0.022 22 96 37.31 0.985 0.020 27 123
Table 18: Per-scene results on TanksAndTemple with Instant-NGP
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Ignatius 28.48 0.947 0.070 63 130 28.41 0.947 0.070 78 137 28.44 0.948 0.069 98 124
Caterpillar 26.30 0.913 0.111 47 178 26.32 0.915 0.109 63 174 26.42 0.915 0.107 92 178
Family 34.33 0.964 0.039 27 152 34.43 0.966 0.037 46 155 34.47 0.967 0.037 69 167
Barn 27.76 0.859 0.205 39 256 28.00 0.865 0.204 57 259 28.10 0.872 0.189 80 250
Truck 27.62 0.912 0.109 49 334 27.85 0.916 0.106 79 349 27.91 0.919 0.104 99 336
avg 28.90 0.919 0.107 45 210 29.00 0.922 0.105 65 215 29.07 0.924 0.101 88 211
Top-1 Ignatius 28.24 0.946 0.077 32 205 28.19 0.928 0.145 48 211 28.46 0.930 0.072 61 209
Caterpillar 26.31 0.913 0.151 27 185 26.60 0.968 0.049 42 187 26.66 0.969 0.145 29 185
Family 34.60 0.965 0.052 25 247 34.66 0.873 0.225 26 252 34.87 0.875 0.049 32 235
Barn 28.31 0.862 0.236 23 232 28.52 0.920 0.128 33 254 28.65 0.920 0.227 22 264
Truck 27.71 0.905 0.136 29 230 27.94 0.949 0.076 46 251 27.96 0.951 0.129 42 251
avg 28.93 0.920 0.131 27 220 29.02 0.924 0.127 39 231 29.16 0.927 0.125 37 229
Top-2 Ignatius 28.24 0.946 0.077 53 381 28.19 0.928 0.145 48 380 28.46 0.930 0.072 59 379
Caterpillar 26.31 0.913 0.151 31 490 26.60 0.968 0.049 42 488 26.66 0.969 0.145 46 480
Family 34.60 0.965 0.052 26 340 34.66 0.873 0.225 26 369 34.87 0.875 0.049 29 379
Barn 28.31 0.862 0.236 23 589 28.52 0.920 0.128 33 604 28.65 0.920 0.227 38 572
Truck 27.71 0.905 0.136 41 690 27.94 0.949 0.076 46 768 27.96 0.951 0.129 65 846
avg 29.04 0.918 0.129 35 498 29.18 0.928 0.125 39 522 29.32 0.929 0.124 47 531
Ens Ignatius 28.17 0.944 0.077 49 362 28.28 0.951 0.076 73 468 28.33 0.950 0.073 90 615
Caterpillar 26.39 0.917 0.147 58 512 26.70 0.931 0.144 67 676 26.87 0.928 0.139 98 876
Family 34.54 0.966 0.053 37 448 34.65 0.970 0.049 52 604 34.89 0.970 0.048 67 776
Barn 28.25 0.865 0.232 27 703 28.48 0.873 0.224 43 900 28.78 0.878 0.219 44 1110
Truck 27.83 0.916 0.134 29 923 27.93 0.921 0.128 40 1275 28.05 0.922 0.124 53 1638
avg 29.25 0.916 0.141 40 590 29.44 0.924 0.137 51 785 29.65 0.924 0.133 66 1003
Table 19: Per-scene results on LLFF with Instant-NGP
Method Scene M=3 M=4 M=5
PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs PSNR SSIM LPIPS $\|w\||_{0}$ GFLOPs
Baseline Fern 23.48 0.741 0.149 95 631 23.20 0.746 0.148 146 676 23.29 0.748 0.143 208 697
Flower 26.94 0.806 0.121 59 611 26.84 0.807 0.120 86 581 26.90 0.809 0.116 112 586
Fortress 27.13 0.821 0.089 44 560 28.24 0.823 0.079 56 447 28.48 0.824 0.075 77 761
Horns 26.00 0.805 0.131 61 472 26.02 0.815 0.123 85 441 26.00 0.818 0.119 114 414
Leaves 17.78 0.555 0.227 120 603 17.60 0.557 0.225 190 614 17.65 0.558 0.223 285 604
Orchids 18.79 0.595 0.186 115 519 18.76 0.558 0.183 182 558 18.64 0.558 0.190 277 564
Room 31.68 0.917 0.084 50 461 31.95 0.919 0.079 65 479 32.02 0.920 0.077 97 476
Trex 26.76 0.876 0.083 35 452 26.83 0.880 0.081 41 438 26.76 0.882 0.080 49 481
avg 24.82 0.764 0.134 72 539 24.93 0.763 0.130 106 529 24.97 0.764 0.128 152 573
Top-1 Fern 23.13 0.729 0.296 55 906 23.55 0.718 0.304 61 893 23.55 0.725 0.301 87 852
Flower 26.49 0.796 0.229 23 786 26.67 0.787 0.227 33 876 26.59 0.788 0.223 44 888
Fortress 26.63 0.800 0.210 23 605 26.73 0.743 0.282 40 1131 27.70 0.798 0.216 37 1141
Horns 25.09 0.765 0.274 37 659 24.90 0.756 0.280 65 745 24.43 0.730 0.311 90 762
Leaves 17.74 0.532 0.386 66 833 17.80 0.532 0.394 104 993 17.78 0.534 0.398 109 937
Orchids 19.09 0.593 0.312 52 654 19.23 0.592 0.319 82 739 19.16 0.590 0.324 89 729
Room 30.63 0.912 0.261 17 987 31.16 0.909 0.252 28 988 31.19 0.908 0.247 31 1083
Trex 25.98 0.867 0.252 28 657 25.62 0.852 0.263 42 691 25.75 0.846 0.262 60 702
avg 24.78 0.762 0.244 37 761 24.88 0.765 24.520 57 882 24.52 0.763 0.239 68 887
Top-2 Fern 23.13 0.729 0.296 82 1422 23.55 0.735 0.280 90 1619 23.46 0.741 0.273 106 1758
Flower 26.49 0.796 0.229 32 1373 26.70 0.800 0.218 36 1360 26.69 0.802 0.212 49 1469
Fortress 26.63 0.800 0.210 26 1086 27.32 0.812 0.190 31 1103 27.82 0.820 0.186 37 1250
Horns 25.09 0.765 0.274 52 1236 25.36 0.780 0.257 60 1176 25.67 0.804 0.240 77 1241
Leaves 17.74 0.532 0.386 104 1484 17.77 0.539 0.370 112 1609 17.87 0.562 0.348 125 1702
Orchids 19.09 0.593 0.312 89 1179 18.95 0.598 0.299 93 1211 19.26 0.618 0.285 117 1309
Room 30.63 0.912 0.261 17 1034 31.16 0.909 0.252 28 988 31.19 0.908 0.247 31 1198
Trex 25.98 0.867 0.252 46 1001 26.07 0.869 0.241 50 1114 26.38 0.880 0.236 56 1197
avg 25.06 0.772 0.241 57 1227 25.11 0.777 0.238 62 1272 25.17 0.777 0.237 75 1391
Ens Fern 23.40 0.748 0.256 87 1839 23.41 0.752 0.252 119 2546 23.43 0.757 0.247 146 3177
Flower 26.92 0.816 0.198 45 1830 27.07 0.822 0.194 76 2447 27.16 0.825 0.190 116 3123
Fortress 28.36 0.830 0.170 49 1765 28.42 0.831 0.169 83 2426 28.55 0.833 0.167 129 2909
Horns 26.23 0.819 0.212 74 1413 26.31 0.827 0.206 88 1913 26.41 0.832 0.201 133 2376
Leaves 17.61 0.550 0.345 109 1862 17.59 0.552 0.339 166 2500 17.57 0.552 0.340 253 3149
Orchids 18.71 0.596 0.288 89 1495 18.72 0.597 0.285 127 2027 18.72 0.599 0.283 197 2583
Room 32.40 0.926 0.229 49 1264 32.61 0.937 0.226 88 1646 32.77 0.935 0.224 108 2004
Trex 26.81 0.890 0.222 35 1390 26.91 0.888 0.221 70 1873 26.90 0.889 0.216 105 2377
avg 25.06 0.772 0.240 67 1607 25.13 0.776 0.237 102 2172 25.19 0.778 0.234 148 2712